### v2

# Federated Learning with Dynamic Client Arrival and Departure: Convergence and Rapid Adaptation via Initial Model Construction

Abstract
Most federated learning (FL) approaches assume a fixed client set. However, real-world scenarios often involve clients dynamically joining or leaving the system based on their needs or interest in specific tasks. This dynamic setting introduces unique challenges: (1) the optimization objective evolves with the active client set, unlike traditional FL with a static objective; and (2) the current global model may no longer serve as an effective initialization for subsequent rounds, potentially hindering adaptation. To address these challenges, we first provide a convergence analysis under a non-convex loss with a dynamic client set, accounting for factors such as gradient noise, local training iterations, and data heterogeneity. Building on this analysis, we propose a model initialization algorithm that enables rapid adaptation to new client sets whenever clients join or leave the system. Our key idea is to compute a weighted average of previous global models, guided by gradient similarity, to prioritize models trained on data distributions that closely align with the current client set, thereby accelerating recovery from distribution shifts. This plug-and-play algorithm is designed to integrate seamlessly with existing FL methods, offering broad applicability in practice. Experimental results on diverse datasets including both image and text domains, varied label distributions, and multiple FL algorithms demonstrate the effectiveness of the proposed approach across a range of scenarios.
1 Introduction
Federated learning (FL) is a decentralized machine learning paradigm that facilitates collaborative model training across multiple clients, such as smartphones and Internet of Things (IoT) clients, without exchanging individual data. Instead of transmitting raw data to the central server, each client performs local training using its proprietary data, sending only model updates to the server. These updates are then aggregated to refine the global model. In conventional FL frameworks, the cohort of clients engaged in training is typically static, implying the objective function is also fixed.
Motivation: However, in practical FL systems, the dynamic nature of client arrival and departure presents significant challenges to maintaining a robust and accurate model. For instance, clients may lose interest when their local data no longer aligns with the central task, such as when a user’s application shifts from text predictions to image recognition. On the other hand, clients may join when their current objectives align with those of other clients, such as when multiple users are working with similar data types or models addressing the same problem, like medical institutions collaborating on disease detection models. Additional factors, such as evolving privacy policies, or shifts in data-sharing preferences, further exacerbate these arrival and departure fluctuations, complicating the overall learning process.
Figure 1: Comparison between traditional FL with partial participation [1, 2, 3, 4, 5, 6] and our proposed dynamic FL framework. Traditional FL approaches assume a fixed client set and a static optimization goal, even when some clients are intermittently available. In contrast, our framework handles dynamic client arrivals and departures, where the optimization goal evolves to minimize the loss for the current client set in each session. A key innovation of our approach is the construction of an adaptive initial model for each session, leveraging a weighted averaging strategy guided by gradient similarity. This initial model prioritizes contributions from clients whose data characteristics align closely with the current set, facilitating efficient adaptation and mitigating the impact of client dynamics. Traditional FL with full client availability is treated as a special case of the static, partial-participation setting in this comparison.
Key Challenges: This dynamic FL setting introduces new challenges that are not present in traditional static FL scenarios with a fixed objective function. Specifically, the arrival and departure of clients dynamically alter the FL system’s objective, as the model must adapt to the current client set and their associated tasks. For instance, if a client contributing unique data classes withdraws, the model no longer needs to classify those classes, fundamentally altering the training objective. Conversely, the addition of clients with previously unrepresented classes necessitates model adaptation to incorporate the updated classification task. Thus, the core challenge is not merely preserving the diversity of the training set but dynamically adjusting the model to align with the evolving tasks defined by the current clients in the system. To address these shifting objectives, FL systems must be designed with the ability to rapidly adapt to the current client set, ensuring the model remains relevant, stable, and effective as the landscape of participating clients evolves.
Illustrative Examples: Figure 1 illustrates our system model, emphasizing its departure from traditional FL with a fixed client set. In conventional FL, although some clients may temporarily become inactive, the overall optimization goal remains static, aiming to minimize the aggregated loss across all clients and converge toward a single global model $\bm{w}^{*}$. Prior works typically focus on designing client sampling strategies (e.g., [6]) or modifying aggregation weights (e.g., [7]) to minimize this static global loss, without accounting for evolving client sets and objectives. In contrast, our approach introduces a session-based framework that inherently supports dynamic client patterns, where the set of active clients evolves across sessions. Instead of assuming a fixed client set throughout training, our system explicitly models client arrivals and departures at the session level. Each session optimizes a session-specific model based on the current client set, ensuring adaptability to shifting data distributions. To enhance efficiency, an initial model is constructed at the start of each session using the last global model in each previous sessions, providing a better starting point than simply resuming from the last model. Within each session, our scheme also supports partial client involvement, allowing training to proceed even when only a subset of clients is active in each round. By jointly addressing dynamic client arrivals across sessions and partial client involvement within sessions, our framework offers a more flexible and realistic approach.
Main Difference from Prior Work: Unlike existing continual learning paradigms that aim to prevent catastrophic forgetting [8, 9, 10], our work focuses on enabling fast adaptation to the current data distribution. This resembles a service provider aiming to build a global model that effectively serves the clients who are currently active, motivated by scenarios (e.g., mobile/edge environments) where clients frequently join and leave. From this perspective, once a client leaves, the priority shifts to optimizing performance for the remaining users rather than preserving accuracy for those who have departed. We thus adopt a dynamic, session-centric view, emphasizing responsiveness to the active client population. In Section 6, we include class incremental learning as one of our baselines and show that its performance is much worse under our objective.
Contributions: Given the challenges introduced by the dynamic nature of client arrival and departure, this work makes the following key contributions to advance the field:
  • We provide a novel theoretical analysis with a non-convex loss function, deriving an upper bound on the gradient norm to quantify performance under any pattern of client arrival and departure in dynamic FL. Our analysis considers a dynamic optimization objective, where each session aims to identify an optimized model that minimizes the loss function of the current clients, contrasting with existing literature that primarily addresses a static goal of minimizing the loss across all clients that have ever participated. The derived bound incorporates several critical factors influencing performance, including: (1) stochastic gradient noise arising from inherent randomness in local updates, (2) the number of local training iterations, which affects convergence behavior, and (3) the non-IIDness of data distribution. This comprehensive analysis offers deeper insights into the trade-offs involved in model training under dynamic client arrivals and departures.
  • Building on our analysis, we propose a plug-and-play algorithm designed to address dynamic optimization goals by constructing an effective initial model for each training session. The algorithm seamlessly integrates with existing FL methods (e.g., FedAvg [11], FedProx [12], SCAFFOLD [13], and FedACG [14] to facilitate rapid adaptation to a new client set in dynamic environments. It initializes the model as a weighted average of previous models, with weights determined by gradient similarity. By leveraging these similarities, the algorithm prioritizes models trained on clients whose data characteristics closely align with those of the current ones, thereby improving adaptation and mitigating performance drop caused by dynamic client activities.
  • Experimental evaluations on diverse datasets spanning both image and text domains demonstrate significant performance gains from our algorithm, particularly in scenarios with sporadic or moderately varying client patterns, highlighting the method’s effectiveness in real-world settings.
2 Related Works
Federated learning (FL) has emerged as a prominent paradigm for distributed machine learning, enabling the collaborative training of models across decentralized data sources while preserving data privacy [11, 15, 12]. One of the foundational works introduced the Federated Averaging (FedAvg) algorithm, which remains a cornerstone of many FL systems [11]. Subsequent studies have explored various aspects of federated learning, including communication efficiency [16], robustness to adversarial attacks [17], and personalization strategies [18]. However, these approaches typically assume a fixed set of participating clients throughout the training process. Reviews of FL techniques often consider static client participation, where all clients are expected to remain available for the entire training duration [15, 12]. This assumption simplifies the modeling of convergence and performance but does not adequately capture real-world scenarios characterized by dynamic client pattern. Addressing the dynamic nature of client arrival and departure is a critical gap in the current literature, motivating the need for adaptive methods that can effectively handle the entry and exit of clients during the training process.
To address the limitations of static client sets in federated learning, research on dynamic client selection and flexible participation has gained momentum, particularly in response to the challenges posed by varying client availability [19, 20, 21, 22, 23, 24, 25, 26, 27, 3]. These studies explore strategies such as optimizing client selection based on resource constraints, modeling participation patterns probabilistically, and employing adaptive algorithms to address the effects of non-IID data. The goal is to enhance overall model performance by strategically managing client participation during training, balancing computational efficiency, communication costs, and data representativeness. However, these works often assume a fixed client set, neglecting the dynamic nature of client arrival and departure. While [28] propose a flexible federated learning framework that allows for inactive clients, incomplete updates, or dynamic client participation, their analysis is limited to scenarios where the optimization goal changes only once, specifically when a single client joins during training. This restricts the applicability of their approach to more complex and realistic patterns of client pattern. Also, their analysis relied on more restrictive assumptions such as strong convexity and universally bounded gradient. There is a line of work in continual FL [29, 30] , where the goal is to build a model that preserves all previously learned knowledge by preventing catastrophic forgetting. In contrast, as previously mentioned, our work focuses on rapidly adapting to the current data distribution to maximize the satisfaction of clients currently active in the system. Given this difference in objectives, our approach significantly outperforms continual FL in scenarios involving dynamic client arrivals and departures, as we will see in Section 6.
3 Dynamic Federated Learning with Client Arrival and Departure
We consider a dynamic FL system where clients may join or leave the system based on their interest or task needs. Clients may join when the model aligns with their objectives, when they have sufficient new data to contribute, or when the system offers financial or computational incentives. Conversely, they may leave if the global model drifts from their needs, if they lack sufficient data, or if the model’s performance is not beneficial. Privacy or security concerns, such as adversarial threats or inadequate privacy guarantees, may also lead clients to leave. Additionally, clients may leave to allocate computational resources to other tasks or due to poor local model validation, while others may rejoin when they see improvements in these factors.
To mathematically characterize these scenarios, each round of FL training, denoted by $g\in\mathcal{G}=\{1,\dots,G\}$, involves a client set collected in $\mathcal{K}^{(g)}=\{1,\dots,K^{(g)}\}$. All clients are connected to a central server, with each client maintaining its own ML models tailored to specific tasks such as pattern recognition and natural language processing. These models are trained locally on client-specific data, allowing the system to leverage diverse data sources while preserving user privacy. At global round $g$, every client $k\in\mathcal{K}^{(g)}$ is linked with a dataset $\mathcal{D}_{k}^{(g)}$, which comprises $D_{k}^{(g)}\triangleq|\mathcal{D}_{k}^{(g)}|$ data samples. Each data sample $d\in\mathcal{D}_{k}^{(g)}$ includes a feature vector $d$ and its corresponding label. A loss function $f(\bm{w},d)$ evaluates the performance of parameter $\bm{w}\in\mathbb{R}^{M}$ where $M$ represents the model’s dimensionality for each data sample $d\in\mathcal{D}_{k}^{(g)}$. For a specific $\bm{w}$, during the $g$-th global round, the local loss at client $k$ is defined as $F_{k}^{(g)}(\bm{w})\triangleq\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{f(\bm{w},d)}{D_{k}^{(g)}}.$ Then, the global loss of the ML model is given by $F^{(k)}(\bm{w})\triangleq\frac{1}{D^{(g)}}\sum_{k\in\mathcal{K}^{(g)}}D_{k}^{(g)}F_{k}^{(g)}(\bm{w})$ where $D^{(g)}\triangleq|\mathcal{D}^{(g)}|$ is the cardinality of $\mathcal{D}^{(g)}$, the collection of data samples across clients. The main objective of training is to minimize the global loss function, which determines the model’s effectiveness in real-world downstream tasks on clients. However, dynamic client pattern, where clients may join or leave the system at any time, causes the global loss function to vary across rounds. Thus, the optimal global model parameters form a sequence $\{\bm{w}^{(g)\ast}\}_{g=1}^{G}$, where
$\bm{w}^{(g)\ast}=\arg\min_{\bm{w}\in\mathbb{R}^{M}}F^{(g)}(\bm{w},\mathcal{D}^{(g)}),\quad g\in\mathcal{G}.$ (1)
Given this dynamic objective, the training process in round $g$ is specifically designed to optimize the model for the data currently held by the participating clients, aiming to minimize the loss function $F^{(g)}(\bm{w},\mathcal{D}^{(g)})$. Unlike continual learning, which typically aims to retain performance on previously encountered data distributions, this training strategy focuses on optimizing for the data distribution specific to the current round $g$. In each training round, the server broadcasts the current global model $\bm{w}^{(g)}\in\mathbb{R}^{M}$ to all clients $\mathcal{K}^{(g)}$. Upon receiving the global model, each client $k\in\mathcal{K}^{(g)}$ updates the model using its local dataset $\mathcal{D}_{k}^{(g)}$ over $e_{k}^{(g)}$ steps of stochastic gradient descent (SGD). During the $h$-th SGD iteration, where $h\in\{0,\dots,e_{k}^{(g)}-1\}$, the update rule is given by $\bm{w}_{k}^{(g),h+1}=\bm{w}_{k}^{(g),h}-\eta^{(g)}\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),h},\mathcal{B}_{k}^{(g)}),$ where $\eta^{(g)}$ denotes the learning rate in the $g$-th round, and $\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),h},\mathcal{B}_{k}^{(g)})$ is the gradient of client $k$’s local loss function evaluated on the mini-batch dataset $\mathcal{B}_{k}^{(g)}\subset\mathcal{D}_{k}^{(g)}$. The mini-batch dataset is sampled from the local dataset $\mathcal{D}_{k}^{(g)}$ to compute the stochastic gradient. The initial model for local training is the current global model, i.e., $\bm{w}_{k}^{(g),0}=\bm{w}^{(g)}$, and the final local model after training is denoted as $\bm{w}_{k}^{(g),\mathsf{F}}$. After completing local training, clients send their final models $\bm{w}_{k}^{(g),\mathsf{F}}$ to the server. The server aggregates these models to update the global model as $\bm{w}^{(g+1)}=\sum_{k\in\mathcal{K}^{(g)}}(D_{k}^{(g)}\bm{w}_{k}^{(g),\mathsf{F}}/D^{(g)}).$ The server initiates the next training round by transmitting the updated global model $\bm{w}^{(g+1)}$ to the new client set $\mathcal{K}^{(g+1)}$. This set includes clients that joined before model aggregation and excludes those that left during the previous round. Unlike prior studies, which optimize for all potential clients, this work focuses on optimizing the model for the current set of participating clients in each round, effectively targeting $\bm{w}^{(g)\ast}$. This strategy aligns with the dynamic nature of client patterns, avoiding optimization for clients that may not rejoin. In this framework, existing clients must complete local training before leaving, while new clients can join anytime and participate in the next global round if they join after the model broadcast.
4 Theoretical Analysis
[…]
Assumption 1 (Smoothness of Loss Functions).
Local loss function $F_{k}^{(g)}$ is $\beta$-smooth, $\forall k\in\mathcal{K}^{(g)},g$: $\|\nabla F_{k}^{(g)}(\bm{w})-\nabla F_{k}^{(g)}(\bm{w}^{\prime})\|\leq\beta\|\bm{w}-\bm{w}^{\prime}\|,~{}\forall\bm{w},\bm{w}^{\prime}\in\mathbb{R}^{M},$ which implies the $\beta$-smoothness of the global loss function $F^{(g)},\;\forall g\in\mathcal{G}$.
Assumption 2 (Bounded Dissimilarity of Local Loss Functions).
There exist finite constants $\zeta_{1}\geq 1$, $\zeta_{2}\geq 0$, for any set of coefficients $\{a_{k}\geq 0\}_{k\in\mathcal{K}^{(g)}}$, where $\sum_{k\in\mathcal{K}^{(g)}}a_{k}=1$, such that $\forall g,\bm{w}$ $\sum\nolimits_{k\in\mathcal{K}^{(g)}}a_{k}\|\nabla F_{k}^{(g)}(\bm{w})\|^{2}\leq\zeta_{1}\|\sum\nolimits_{k\in\mathcal{K}^{(g)}}a_{k}\nabla F_{k}^{(g)}(\bm{w})\|^{2}+\zeta_{2}$.
The parameters $\zeta_{1}$ and $\zeta_{2}$, introduced in Assumption 2, quantify the degree of heterogeneity (i.e., the extent of non-i.i.d. behavior) among the datasets distributed across devices. When the datasets are fully homogeneous (i.e., i.i.d. across devices), these parameters take the values $\zeta_{1}=1$ and $\zeta_{2}=0$. However, as the heterogeneity of the data increases, both $\zeta_{1}$ and $\zeta_{2}$ will grow accordingly. These parameters play crucial roles in our convergence analysis, as they influence the allowable step size for SGD and the overall convergence rate of ML model training.
To further capture the heterogeneity within the local datasets of individual devices, we introduce the concept of local data variability.
Definition 1 (Local Data Variability).
The local data variability at each device $k$ is measured via $\Theta_{k}\geq 0$, which $\forall\bm{w},k$ satisfies $\|\nabla f(\bm{w},d)-\nabla f(\bm{w},d^{\prime})\|\leq\Theta_{k}^{(g)}\|d-d^{\prime}\|,\quad\forall d,d^{\prime}\in\mathcal{D}_{k}^{(g)}$. We further define $\Theta^{(g)}=\max_{k\in\mathcal{K}^{(g)}}\{\Theta_{k}^{(g)}\}$.
We next obtain the convergence analysis for the non-convex loss function in the dynamic system.
Theorem 1 (Upper Bound on the Gradient Norm).
Suppose each client performs local SGD updates and the server aggregates via weighted averaging. Define $e_{\text{max}}^{(g)}=\max_{k\in\mathcal{K}^{(g)}}e_{k}^{(g)}$ and $e_{\text{min}}^{(g)}=\min_{k\in\mathcal{K}^{(g)}}e_{k}^{(g)}$. Assuming the learning rate $\eta^{(g)}$ is less than
$\min\!\Bigl{\{}(0.5-((e_{\text{min}}^{(g)})^{-1}-1)^{2})\,\beta^{-1},\;\sqrt{\Lambda^{(g)}}\,(2\beta\sqrt{(\zeta_{1}+\Lambda^{(g)})\,e_{\text{max}}^{(g)}(e_{\text{max}}^{(g)}-1)})^{-1}\Bigr{\}},$
where $\Lambda^{(g)}<1$ is a constant, then for any pattern of client arrivals and departures, the cumulative average of the gradient of the global loss over $G$ rounds satisfies:
$\frac{1}{G}\sum_{g=1}^{G}\mathbb{E}\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}\leq\underbrace{\frac{1}{G}\sum_{g=1}^{G}\frac{2\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)})-F^{(g)}(\bm{w}^{(g+1)})\bigr{]}}{\eta^{(g)}(1-\Lambda^{(g)})}}_{\text{(a)}}+\underbrace{\frac{1}{G}\sum_{g=1}^{G}\frac{4\beta(\Theta^{(g)})^{2}\eta^{(g)}}{1-\Lambda^{(g)}}\sum_{k\in\mathcal{K}^{(g)}}\frac{(\widetilde{\sigma}_{k}^{(g)})^{2}\,e_{k}^{(g)}}{B_{k}^{(g)}}\frac{D_{k}^{(g)}-B_{k}^{(g)}}{(D_{k}^{(g)})^{2}}}_{\text{(b)}}$
$\quad+\underbrace{\frac{1}{G}\Bigl{(}\sum_{g=1}^{G}\frac{4\beta^{2}(\eta^{(g)})^{2}}{D^{(g)}(1-\Lambda^{(g)})}\Bigr{)}\sum_{k\in\mathcal{K}^{(g)}}(e_{k}^{(g)}-1)\Bigl{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigr{)}\frac{\Theta^{(g)}(\widetilde{\sigma}_{k}^{(g)})^{2}}{B_{k}^{(g)}}}_{(c)}+\underbrace{\frac{1}{G}\sum_{g=1}^{G}8\beta^{2}(\eta^{(g)})^{2}(e_{\max}^{(g)})(e_{\max}^{(g)}-1)\frac{\zeta_{2}}{1-\Lambda^{(g)}}}_{\text{(d)}}.$
Proof.
Please see Appendix E. ∎
Term (a) reflects the expected improvement in the loss function at each training stage. Terms (b) and (c) capture the negative impacts stemming from mini-batch sampling noise (controlled by batch size $B_{k}^{(g)}$), local data heterogeneity across clients (measured by $\Theta^{(g)}$), and the number of local SGD steps ($e_{k}^{(g)}$). Term (d) accounts for the influence of the maximum number of local SGD steps ($e_{\text{max}}^{(g)}$) and inter-client model dissimilarity (represented by $\zeta_{2}$).
Inspiration from Theorem 1 for Initial Model Construction. Term (a) in the bound (2) decomposes neatly into two contributions:
$\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)})-F^{(g)}(\bm{w}^{(g)\ast})\bigr{]}\quad\text{and}\quad\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)\ast})-F^{(g)}(\bm{w}^{(g+1)})\bigr{]}.$
The second of these arises from the update that follows round $g$ and is beyond our control at that stage. Our algorithmic leverage lies in the first term: by constructing an initial model $\bm{w}_{\mathrm{init}}^{(g)}$ aiming to satisfy
$\mathbb{E}\bigl{[}F^{(g)}(\bm{w}_{\mathrm{init}}^{(g)})-F^{(g)}(\bm{w}^{(g)\ast})\bigr{]}\;\leq\;\mathbb{E}\bigl{[}F^{(g)}(\bm{w}^{(g)})-F^{(g)}(\bm{w}^{(g)\ast})\bigr{]},$
we directly tighten the upper bound in Theorem 1 and reduce the cumulative average gradient norm.
Algorithmic Realization via Pilot-Model Similarity Weighting. When the data distribution in round $g$ closely matches that of a previous session, reusing that session’s final model as $\bm{w}_{\mathrm{init}}^{(g)}$ naturally yields the desired reduction in Term (a). In more heterogeneous or evolving settings, however, no single past model suffices. Instead, we form
$\bm{w}_{\mathrm{init}}^{(g)}=\sum_{s<g}\alpha_{s}\,\bm{w}^{(s)},$
where the weights $\{\alpha_{s}\}$ are determined by the similarity between each prior data distribution and the current one. We estimate similarity by computing pilot-model gradients: greater alignment of gradients indicates higher relevance and thus a larger $\alpha_{s}$. This similarity-weighted ensemble initialization both accelerates convergence in stationary regimes and provides robust adaptation under distribution shift, directly minimizing the controllable component of our convergence bound. From this argument, it follows that our similarity-weighted warm-start strategy is agnostic to the choice of federated learning algorithm and can be seamlessly integrated into any aggregation scheme; we validate this generality by experiments in Section 6. To make this concrete, in the next section we describe how to construct the similarity weights $\{\alpha_{s}\}$ that underpin our warm-start strategy.
5 Dynamic Initial Model Construction for Fast Adaptation
Terminology: Algorithm 1 outlines our proposed “Dynamic Initial Model Construction” algorithm. In this context, a “session” differs from the “number of global rounds.” A new session is initiated whenever there is a change in the data distribution, such as clients joining or leaving, while the client set remains constant within a session. Each session comprises at least one global round.
Training: From 8 to 11, each client in the current set performs local model training based on either a new initial model (9) or the latest model at the server (10). Upon completion of local training, each client transmits its final local model to the server, where global aggregation is performed. It is worth noting that in Algorithm 1, the local model training and the global aggregation can be performed in any desired way (e.g., FedAvg, FedProx, SCAFFOLD, and FedACG). In 12, we only save the last global model in each session, either for pilot model computation or as components of a new initial model.
Algorithm 1 Dynamic Initial Model Construction
$T$: global rounds per session, $P$: pilot preparation sessions, $S$: total sessions $(S\geq P)$, $V$: gradient computation global rounds.
[…]
3:Local training starting from $\bm{w}_{s}^{(v)}$ for $e_{k}^{(sT+t)}$ steps.
[…]
Pilot Model: As defined in 13, the pilot model is established upon the completion of the pilot preparation stage (the first $P$ sessions). It is calculated as the simple average of the final global model from each of these $P$ sessions:
$\bm{w}_{\text{Pilot}}=\frac{\sum_{s=0}^{P-1}\bm{w}^{(sT+T-1)}}{P}\triangleq\frac{\sum_{s=0}^{P-1}\bm{w}_{s}^{\text{last}}}{P}$ (3)
We introduce the notation $\bm{w}_{s}^{\text{last}}\triangleq\bm{w}^{(sT+T-1)}$ to concisely represent the global model saved at the end of session $s$. These $P$ saved global models are subsequently discarded to free memory.
Gradient Computation and Initial Model Construction: Following the pilot model preparation stage, each subsequent session $s$ starts with a gradient computation process designed to capture the underlying data distribution. This process uses the pilot model as the initial model, $\bm{w}_{s}^{(0)}=\bm{w}_{\text{Pilot}}$, and runs V additional global rounds to obtain a final model. $\bm{w}_{s}^{(V-1)}$ (2 to 5). The computed gradient $G_{s}$ is then defined as the difference $\bm{w}_{s}^{(V-1)}-\bm{w}_{\text{Pilot}}$ (7):
$G_{s}\triangleq\bm{w}_{s}^{(V-1)}-\bm{w}_{s}^{(0)}=\bm{w}_{s}^{(V-1)}-\bm{w}_{\text{Pilot}}.$ (4)
This quantity is particularly useful for evaluating the similarity between different data distributions, and its value is independent of client data amount. For each session following the pilot preparation stage, only one computed gradient is retained. If at least one such gradient has been saved after the pilot stage (6), the initial model for session $s$ is constructed using only the models saved after the pilot stage.
$\bm{w}_{s}^{\text{init}}\triangleq\sum\nolimits_{z=P}^{s-1}\alpha_{s,z}\bm{w}_{z}^{\text{last}},\quad\alpha_{s,z}=\frac{e^{-R\|G_{s}-G_{z}\|_{2}}}{\sum_{z=P}^{s-1}e^{-R\|G_{s}-G_{z}\|_{2}}}.$ (5)
The update rule in (5) defines a weighted average of the saved models $\bm{w}_{z}^{\text{last}}$, where $\alpha_{s,z}$ denotes the corresponding weight. A small difference between the past computed gradient $G_{z}$ and the current gradient $G_{s}$ suggests that the data distribution in the $z$-th session is similar to that of the current session. Since the goal of each session is to learn a model that performs well on the current data distribution, the corresponding saved model $\bm{w}_{z}^{\text{last}}$ serves as a suitable initialization. In such cases, a smaller $\|G_{s}-G_{z}\|_{2}$ yields a larger $\alpha_{s,z}$, thereby assigning more weight to $\bm{w}_{z}^{\text{last}}$ in (5) and making it a dominant component of the constructed initial model.
The constant $R$ in $\alpha_{s,z}$ controls the sensitivity to gradient differences. When $R=0$, the initial model reduces to the simple average of all saved models after the pilot preparation stage, regardless of data similarity. In contrast, as $R\rightarrow\infty$, the initial model converges to the saved model associated with the most similar data distribution.
Remark 1.
Our dynamic initial model construction algorithm provides a simple, effective, and easily integrated approach compatible with any federated learning method. Although providing formal guarantees for warm-start strategies is notoriously challenging [33, 34], especially under client heterogeneity and churn, preliminary theoretical insights (Theorem 1) and extensive empirical validation in Section 6 demonstrate its practical utility. To keep the main text focused, we defer detailed discussions to the appendix: Appendix B covers gradient similarity, sensitivity experiments on $R$ and how well computed gradients capture the underlying data distribution, as well as ablation studies; Appendix C analyzes computation, communication and storage costs.
6 Experiments
FL Algorithm and Baselines: We evaluate the proposed initial model construction method by integrating it with established FL algorithms, including FedAvg [11], FedProx [12], SCAFFOLD [13] and FedACG [14]. Notably, FedACG represents a state-of-the-art approach that has demonstrated superior performance, surpassing methods such as FedDyn [35]. For comparative analysis, we also consider three baseline approaches: Previous, Average, and Continuous (for image dataset). The Previous baseline utilizes the model from the preceding session for continued training, foregoing the explicit construction of a dedicated initial model [36, 37, 11, 38]. The Average baseline initializes models using the unweighted average of prior global models. The Continuous baseline employs the Class-Semantic Relation Distillation Loss [29], which uses the model from the previous session as an anchor for the current client local model. This encourages the new model to preserve relationships between class representations from prior sessions while adapting to incoming client data.
Table 1: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and FedACG under different label distributions for CIFAR datasets.
Dataset Overlap Dir($\alpha$) FL Alg. First Session Transition Second Session Transition
Proposed Continuous Previous Average Proposed Continuous Previous Average
CIFAR10 0.0 0.3 FedAvg 65.52±2.77 68.96±2.12 69.60±2.51 74.63±1.67 63.85±5.30 61.25±5.54 47.61±1.54 64.37±4.06
FedProx 69.88±3.95 69.12±2.22 71.43±2.77 74.63±2.13 68.07±3.28 61.19±5.87 52.76±2.63 67.66±2.70
SCAFFOLD 80.49±1.08 76.39±2.37 74.21±1.45 80.13±1.31 77.25±1.05 68.60±4.61 72.88±0.83 72.79±1.11
FedACG 80.10±3.69 69.03±2.41 35.46±7.74 51.24±8.33 73.35±4.07 61.65±5.26 32.21±2.56 56.68±6.25
0.7 FedAvg 79.15±1.31 77.33±0.66 76.04±2.09 79.91±0.36 74.63±2.55 70.22±2.81 50.92±1.53 70.63±1.89
FedProx 79.32±1.32 77.35±0.58 77.70±0.48 80.77±0.55 76.29±0.93 70.50±2.54 56.36±0.81 72.05±1.44
SCAFFOLD 84.62±0.36 82.96±0.03 79.59±0.80 83.43±0.38 81.15±0.90 71.46±5.91 75.57±2.02 74.88±1.05
FedACG 86.33±1.22 77.49±0.70 40.10±5.43 62.38±9.57 82.04±2.98 70.40±2.41 36.22±1.92 58.53±9.42
0.2 0.3 FedAvg 61.96±7.12 63.19±3.24 67.95±2.66 69.11±1.35 62.81±5.75 61.23±3.19 57.71±3.46 61.58±1.24
FedProx 67.42±5.09 63.69±3.15 68.76±1.77 68.99±1.37 67.07±3.80 60.94±4.36 60.95±4.89 64.41±2.09
SCAFFOLD 76.45±1.20 70.17±4.42 69.76±2.06 76.31±2.25 77.39±0.94 67.34±3.10 71.34±3.17 72.76±0.35
FedACG 74.96±2.18 63.57±3.40 39.80±2.28 57.44±7.00 76.82±2.60 59.94±4.85 36.45±2.19 59.72±3.47
0.7 FedAvg 74.17±3.36 71.04±0.62 73.42±0.51 74.22±0.82 73.14±4.08 70.57±1.29 64.26±0.87 67.37±0.63
FedProx 75.71±0.30 70.91±0.67 73.83±0.61 73.76±1.28 76.05±1.53 70.43±1.41 66.23±1.13 69.14±1.65
SCAFFOLD 80.21±0.38 76.19±0.68 75.92±0.42 79.58±0.62 80.75±1.08 73.19±3.29 76.24±1.06 76.29±1.36
FedACG 80.68±0.40 70.99±0.54 47.62±8.73 71.47±8.01 81.66±0.58 70.52±1.35 38.88±5.44 71.23±2.63
CIFAR 100 0.0 0.3 FedAvg 47.87±1.26 41.71±1.20 44.60±0.91 47.33±1.09 51.33±0.82 42.88±1.59 45.39±1.93 41.50±3.71
FedProx 47.09±0.54 41.89±0.93 45.80±1.23 47.73±0.18 50.92±0.53 43.20±1.37 48.81±0.84 44.11±1.10
SCAFFOLD 51.68±1.00 44.51±1.14 47.77±1.35 52.35±0.64 54.64±1.57 43.34±0.45 50.74±2.09 49.16±1.10
FedACG 45.83±1.52 41.61±0.32 3.82±2.15 4.01±1.80 48.33±5.25 46.63±1.27 2.00±0.40 10.67±8.36
0.7 FedAvg 52.14±0.29 46.18±0.56 47.83±0.25 51.04±0.59 54.60±0.87 45.78±1.61 48.30±0.75 46.65±1.41
FedProx 50.77±0.70 46.17±1.30 49.77±0.43 51.32±0.64 54.25±0.69 46.56±0.26 52.34±0.72 47.37±0.99
SCAFFOLD 55.00±0.09 52.25±0.79 52.01±0.88 55.66±0.55 57.60±0.27 48.93±1.36 54.52±0.91 51.63±1.06
FedACG 52.97±4.84 46.29±1.05 1.82±1.49 5.91±6.76 55.38±6.27 50.71±0.46 2.59±1.47 22.75±15.49
0.2 0.3 FedAvg 46.95±1.24 42.03±0.28 44.78±0.61 46.11±0.63 50.29±1.53 43.41±0.94 47.26±2.04 42.89±0.67
FedProx 46.35±0.51 42.21±0.16 45.08±0.73 47.24±0.48 50.61±0.16 43.74±1.19 48.53±0.80 44.06±0.20
SCAFFOLD 50.20±1.16 45.94±0.79 48.23±0.49 51.01±0.84 54.04±0.73 45.23±0.73 52.18±0.84 49.08±0.27
FedACG 56.46±0.94 42.83±0.70 20.87±0.28 34.79±5.97 57.58±0.28 44.27±0.33 22.31±2.51 41.11±4.48
0.7 FedAvg 50.56±0.28 46.62±0.23 48.28±0.52 48.64±0.52 54.05±0.70 48.38±0.15 52.18±0.13 45.85±1.05
FedProx 48.90±0.57 46.60±0.41 48.33±1.52 50.24±0.42 53.12±0.82 48.24±0.41 52.41±0.83 47.99±0.95
SCAFFOLD 53.27±0.68 52.21±0.25 50.87±0.91 54.09±0.21 56.76±0.77 51.71±0.42 55.09±0.65 51.86±0.28
FedACG 59.11±0.72 46.66±0.40 23.82±0.84 35.15±9.77 60.70±0.39 48.22±0.06 24.38±1.71 47.24±0.42
Task and Dataset: We perform classification tasks on seven diverse benchmarks, including six image datasets (MNIST [39], Fashion-MNIST [40], SVHN [41], CIFAR-10 [42], CIFAR-100 [42], TinyImageNet [42]) and the AG News text corpus [43]. Results for CIFAR-10, CIFAR-100, and AG News are shown in the main text. Additional experiments appear in Appendix A.
Label Distribution: In our main experiments, we utilize 20 clients, while results with more clients (e.g. 50 or 100) are provided in Appendix A and the supplementary materials. Data samples are distributed among clients by defining the set of labels available within a training session. To introduce variation in the label distribution across the training process, the set of labels available in one training session can partially overlap with the set of labels available in a subsequent session. This overlap level can range from zero (meaning the label sets for consecutive sessions are disjoint) up to a value approaching one (meaning most labels are shared, with only a small fraction unique to each set). For the labels available in a session, they are distributed among the active clients using a Dirichlet distribution parameterized by $\alpha$ [11, 44, 45, 46, 47]. A smaller value of $\alpha$ results in a more heterogeneous distribution of labels across clients, leading to higher non-IIDness between clients. Conversely, a larger value of $\alpha$ yields a more homogeneous distribution of labels across clients.
Table 2: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and FedACG under different label distributions for AG News text corpus.
Dataset Overlap Dir($\alpha$) FL Alg. First Session Transition Second Session Transition
Proposed Previous Average Proposed Previous Average
AG News 0.0 0.3 FedAvg 68.34±0.34 14.96±0.17 68.94±0.45 73.96±0.84 14.09±0.30 14.30±0.04
FedProx 68.29±0.37 14.95±0.17 68.90±0.46 73.93±0.85 14.08±0.31 14.29±0.03
SCAFFOLD 61.84±0.45 13.13±0.44 60.85±2.03 66.68±1.80 12.96±0.12 12.96±0.34
FedACG 74.71±0.23 16.16±0.21 58.95±1.50 79.91±0.20 15.14±0.10 26.21±1.36
0.7 FedAvg 68.37±0.38 14.95±0.17 68.96±0.48 73.93±0.81 14.11±0.31 14.30±0.03
FedProx 68.32±0.38 14.95±0.16 68.92±0.46 73.89±0.86 14.11±0.31 14.28±0.04
SCAFFOLD 61.90±0.45 13.14±0.43 60.90±2.19 66.73±1.86 12.97±0.14 12.98±0.34
FedACG 74.74±0.21 16.16±0.21 58.96±1.57 79.88±0.25 15.14±0.11 26.18±1.36
0.6 0.3 FedAvg 57.94±0.93 52.12±1.33 51.27±0.93 60.53±0.87 52.54±1.24 52.26±0.91
FedProx 57.86±0.93 52.10±1.33 51.22±0.90 60.45±0.91 52.50±1.24 52.23±0.91
SCAFFOLD 47.88±1.57 47.65±1.57 45.27±1.25 49.51±1.06 49.29±1.49 46.74±1.09
FedACG 67.24±0.91 55.26±0.72 57.44±0.76 70.28±0.88 55.11±1.10 57.69±0.68
0.7 FedAvg 57.88±0.93 52.11±1.27 51.26±0.91 60.50±0.93 52.53±1.25 52.28±0.92
FedProx 57.82±0.95 52.09±1.28 51.22±0.88 60.45±0.94 52.51±1.26 52.25±0.92
SCAFFOLD 48.10±1.49 47.73±1.56 45.34±1.24 49.59±1.07 49.33±1.41 46.83±1.13
FedACG 67.24±0.87 55.21±0.72 57.45±0.76 70.34±0.95 55.10±1.12 57.69±0.73
Results and Takeaways: All values are reported as mean $\pm$ sample standard deviation (%) over three random seeds. Table 1 and Table 2 compare the mean accuracy ($\pm$ std) of our proposed warm start algorithm combined with several federated learning methods against the baseline aggregation schemes in the initial rounds following each session transition on CIFAR10, CIFAR100 and AG News. Across diverse datasets and label distribution scenarios, our approach consistently mitigates the accuracy drop induced by dynamic client arrivals and departures and accelerates recovery whenever the new distribution closely resembles one encountered previously. Figure 2 shows the performance of our proposed algorithm for selected setups across different datasets. It visualizes the fast recovery and robustness client arrivals and departures of our proposed algorithm.
Additional Results: We additionally evaluate our approach with a larger number of clients (e.g., 50 or 100) and on more complex datasets such as TinyImageNet in Appendix A and supplementary materials to further demonstrate its effectiveness. We also examine different label distributions and analyze the impact of the scaling factor $R$ in the appendix and supplementary material.
Figure 2: Performance comparison of the proposed algorithm implemented with selected FL algorithms across selected datasets. The results demonstrate the robustness of our proposed scheme to dynamic data distributions caused by client arrivals and departures.
7 Conclusion
This work addresses the challenges posed by dynamic client participation in FL, where clients may join or leave the system during training. We conduct a theoretical analysis under non-convex loss settings, revealing how factors such as stochastic gradient noise, local training iterations, and non-IID data distributions influence performance. Building on this analysis, we propose a plug-and-play algorithm that constructs an effective initial model for each training round using a weighted averaging strategy guided by gradient similarity. The approach integrates seamlessly with existing FL methods, prioritizing models from clients with data characteristics similar to the active clients, thereby enhancing adaptability and mitigating performance degradation. Experimental results demonstrate significant improvements across diverse datasets and client arrival and departure patterns, underscoring the practical utility of the method in real-world FL scenarios. Future work could refine the initial model construction process to trigger updates only when necessary, reducing computational overhead while maintaining performance and enabling more efficient, adaptive FL systems.
Appendix A Extended Results
A.1 Supplementary Results on Label Distributions from the Main Text
A.1.1 20 clients
Table 3 further illustrates the effectiveness of Algorithm 1 on MNIST, Fashion-MNIST and SVHN dataset for a system of 20 clients.
Table 4 further illustrates the effectiveness of Algorithm 1 on TinyImagenet dataset for a system of 20 clients.
Table 3: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and FedACG under different label distributions for MNIST, Fashion-MNIST (FMNIST), and SVHN dataset for 20 clients.
Dataset Overlap Dirichlet
$\alpha$
FL Alg. First Session Transition Second Session Transition
Proposed Continuous Previous Average Proposed Continuous Previous Average
MNIST 0.0 0.3 FedAvg 90.97±0.16 86.47±0.32 62.34±0.64 90.47±0.26 94.06±0.30 84.33±0.58 72.21±0.16 82.05±0.15
FedProx 90.96±0.16 86.47±0.32 62.27±0.62 90.47±0.26 94.06±0.30 84.33±0.58 72.16±0.16 82.01±0.17
SCAFFOLD 88.29±0.42 79.20±7.56 39.90±1.49 86.68±0.45 92.73±0.69 80.35±3.54 53.76±0.41 61.38±0.93
FedACG 93.57±0.12 86.47±0.32 27.49±0.35 74.72±0.69 95.85±0.11 84.33±0.58 26.43±0.17 90.49±0.18
0.7 FedAvg 91.01±0.13 86.98±0.21 63.11±0.78 90.54±0.24 94.09±0.28 84.61±0.52 72.59±0.02 82.33±0.18
FedProx 91.01±0.13 86.98±0.21 63.04±0.78 90.53±0.25 94.09±0.29 84.61±0.52 72.55±0.01 82.30±0.18
SCAFFOLD 88.60±1.34 87.79±2.04 40.67±1.18 87.71±0.84 93.20±0.29 86.64±1.44 53.38±0.77 61.22±1.60
FedACG 93.55±0.08 86.98±0.21 27.53±0.32 74.77±0.61 95.85±0.13 84.61±0.52 26.44±0.19 90.52±0.21
0.2 0.3 FedAvg 92.08±0.03 87.97±0.35 69.69±0.43 90.50±0.10 94.38±0.31 86.46±0.57 80.23±0.35 85.26±0.48
FedProx 92.07±0.03 87.97±0.35 69.66±0.43 90.49±0.10 94.37±0.30 86.46±0.57 80.21±0.35 85.23±0.48
SCAFFOLD 90.30±0.15 85.52±2.54 52.62±1.92 86.13±0.91 93.27±0.66 86.19±1.87 66.33±0.28 68.84±2.03
FedACG 94.62±0.02 87.97±0.35 45.28±0.47 79.06±0.14 96.15±0.09 86.46±0.57 49.42±0.07 91.43±0.08
0.7 FedAvg 92.13±0.06 88.28±0.34 70.44±0.44 90.62±0.08 94.41±0.27 86.74±0.52 80.53±0.28 85.48±0.44
FedProx 92.13±0.06 88.28±0.34 70.38±0.45 90.61±0.08 94.41±0.27 86.74±0.52 80.50±0.25 85.45±0.42
SCAFFOLD 89.66±0.39 89.91±0.67 54.13±0.68 85.91±1.16 93.55±0.24 90.48±1.29 66.33±0.36 68.69±2.34
FedACG 94.63±0.00 88.28±0.34 45.43±0.37 79.18±0.17 96.14±0.07 86.74±0.52 49.39±0.11 91.44±0.02
FMNIST 0.0 0.3 FedAvg 88.16±0.13 83.95±0.54 79.52±0.61 87.95±0.26 85.06±0.12 46.06±5.63 82.70±0.18 81.63±0.66
FedProx 88.14±0.14 83.95±0.54 79.49±0.59 87.93±0.28 85.06±0.13 46.06±5.63 82.65±0.18 81.59±0.67
SCAFFOLD 83.94±0.07 65.52±2.17 64.28±1.37 82.68±0.34 80.05±0.33 67.19±2.81 53.47±0.19 58.81±0.55
FedACG 91.34±0.08 83.95±0.54 65.95±0.34 75.85±0.26 87.75±0.14 46.06±5.63 52.30±0.79 84.53±0.17
0.7 FedAvg 88.24±0.11 84.62±0.26 79.96±0.27 88.10±0.27 85.19±0.11 54.59±4.77 82.83±0.17 81.77±0.54
FedProx 88.23±0.12 84.62±0.26 79.94±0.27 88.08±0.27 85.18±0.10 54.59±4.77 82.81±0.17 81.75±0.56
SCAFFOLD 83.92±0.18 79.86±0.28 64.43±0.36 82.31±0.79 80.34±0.49 72.03±3.14 52.59±0.55 58.90±1.10
FedACG 91.36±0.13 84.62±0.26 65.93±0.43 75.57±0.56 87.88±0.11 54.59±4.77 52.54±0.58 84.41±0.27
0.2 0.3 FedAvg 85.50±0.13 82.21±0.99 79.51±0.16 85.62±0.14 84.73±0.07 70.83±0.96 81.61±0.34 79.63±0.45
FedProx 85.48±0.14 82.21±0.99 79.45±0.14 85.59±0.12 84.72±0.08 70.83±0.96 81.55±0.34 79.65±0.46
SCAFFOLD 80.26±0.62 67.82±5.57 67.80±0.96 76.13±0.11 79.50±0.45 71.21±6.86 58.52±0.91 57.96±2.22
FedACG 88.68±0.03 82.21±0.99 67.71±1.23 78.04±0.86 86.69±1.01 70.83±0.96 61.93±0.58 81.48±0.17
0.7 FedAvg 85.67±0.23 82.90±0.24 79.81±0.38 85.67±0.25 84.86±0.05 73.12±1.04 81.81±0.20 79.27±0.33
FedProx 85.65±0.23 82.90±0.24 79.75±0.39 85.67±0.22 84.85±0.05 73.12±1.04 81.77±0.19 79.32±0.32
SCAFFOLD 80.79±0.40 78.60±2.68 67.97±0.77 76.83±0.17 80.07±0.33 75.83±3.23 57.41±0.31 56.79±2.21
FedACG 88.72±0.06 82.90±0.24 67.17±1.28 77.95±0.69 87.03±0.31 73.12±1.04 61.93±0.46 81.68±0.10
SVHN 0.0 0.3 FedAvg 89.43±0.43 80.34±2.95 84.37±0.36 89.46±0.92 92.69±0.08 87.43±1.24 84.04±1.15 88.79±1.74
FedProx 89.39±0.44 80.20±3.17 84.44±0.45 89.51±0.84 92.71±0.08 87.48±1.16 84.48±1.30 89.22±1.18
SCAFFOLD 65.52±9.48 34.74±5.37 40.04±5.91 56.61±8.63 90.07±1.02 63.16±12.07 60.46±5.96 58.08±13.35
FedACG 92.38±0.39 80.26±3.05 26.38±5.81 26.80±5.87 93.56±0.70 87.43±1.21 25.13±15.30 56.32±18.88
0.7 FedAvg 89.93±0.26 83.57±1.11 85.14±0.92 89.64±0.29 92.98±0.15 88.14±0.89 83.14±3.27 87.42±3.18
FedProx 89.84±0.33 83.60±1.11 85.27±0.92 89.71±0.29 92.97±0.17 88.12±0.87 83.68±3.74 87.54±3.35
SCAFFOLD 79.69±0.25 51.29±6.26 47.24±6.15 72.90±7.94 91.57±0.47 79.18±2.81 71.73±5.85 75.57±4.24
FedACG 92.65±0.22 83.54±1.18 22.31±3.01 29.67±9.71 93.00±0.33 88.11±0.89 23.15±8.56 54.31±14.24
0.2 0.3 FedAvg 90.33±0.43 83.84±2.31 84.83±1.52 89.32±0.59 93.07±0.06 88.46±0.71 87.80±1.12 86.26±2.24
FedProx 90.33±0.44 83.73±2.43 85.08±1.40 89.42±0.61 93.07±0.05 88.47±0.69 88.40±1.03 86.51±2.08
SCAFFOLD 75.64±4.35 44.25±1.33 56.22±3.60 72.17±1.52 89.71±0.89 76.42±5.31 76.17±4.81 74.14±6.65
FedACG 93.00±0.64 83.74±2.28 37.38±1.85 39.66±2.93 91.35±4.27 88.47±0.68 32.39±1.71 66.63±2.01
0.7 FedAvg 90.68±0.23 86.18±0.75 84.57±0.81 89.70±0.42 93.31±0.12 89.13±0.60 87.28±2.73 86.36±4.72
FedProx 90.69±0.23 86.16±0.79 84.94±0.82 89.81±0.38 93.28±0.10 89.12±0.60 87.68±2.56 86.48±4.78
SCAFFOLD 79.47±2.44 57.46±9.02 59.95±4.09 70.40±0.72 91.53±0.55 82.85±1.98 80.87±2.55 81.66±2.25
FedACG 93.42±0.29 86.17±0.80 34.40±1.56 41.88±3.73 94.13±0.32 89.09±0.60 33.71±4.84 62.30±9.89
Table 4: Performance comparison of implementing Algorithm 1 with MOON, FedProx, and FedACG under different label distributions for TinyImagenet dataset for 20 clients.
Dataset Overlap Dirichlet
$\alpha$
FL Alg. First Session Transition Second Session Transition
Proposed Continuous Previous Average Proposed Continuous Previous Average
TinyImagenet 0.0 0.3 MOON 62.93±0.46 53.91±0.94 52.49±1.77 61.26±0.76 61.22±2.52 55.30±0.69 56.06±0.47 57.51±1.15
FedProx 62.01±0.32 53.78±1.00 58.90±0.38 61.69±0.44 60.91±1.42 54.74±0.72 62.27±0.71 57.23±1.37
FedACG 64.08±0.49 53.16±0.79 35.49±8.94 54.57±0.74 62.93±0.57 54.91±0.49 21.86±4.62 48.63±1.41
0.7 MOON 64.29±0.58 56.95±0.49 55.85±0.65 63.36±0.46 63.71±0.62 57.02±0.49 57.24±0.62 60.87±0.36
FedProx 63.32±0.33 57.04±0.56 60.80±0.38 62.85±0.74 63.09±0.63 56.94±0.38 63.80±0.37 61.17±0.60
FedACG 65.59±0.46 56.60±0.56 41.68±2.37 57.52±1.94 63.77±1.61 56.73±0.27 28.79±1.08 55.93±2.33
0.2 0.3 MOON 61.69±0.68 55.14±0.76 55.63±0.84 60.71±0.75 60.20±1.59 57.48±0.64 59.48±0.55 59.13±1.08
FedProx 60.30±0.82 55.26±0.95 57.99±0.54 60.81±0.36 59.29±2.37 57.63±0.35 62.54±0.30 58.28±0.78
FedACG 63.20±0.49 55.40±0.77 51.86±0.54 61.09±0.57 61.78±0.88 57.42±0.62 55.46±0.62 55.05±2.20
0.7 MOON 63.12±0.44 57.12±0.68 58.69±0.63 62.18±0.80 61.83±1.24 59.22±0.82 61.60±0.43 61.45±0.51
FedProx 61.83±0.64 57.14±0.98 58.86±0.39 62.16±0.69 60.86±0.93 58.96±0.62 63.86±0.57 60.70±0.51
FedACG 64.54±0.27 57.46±0.42 54.03±0.98 62.27±0.41 64.17±1.55 59.23±0.55 56.70±0.48 59.89±0.67
A.1.2 100 clients
Table 5 further illustrates the effectiveness of Algorithm 1 on MNIST, Fashion-MNIST, SVHN, CIFAR-10, and CIFAR-100 dataset for a system of 100 clients.
Table 5: Performance comparison of implementing Algorithm 1 with MOON, FedProx, SCAFFOLD and FedACG under different label distributions for MNIST, Fashion-MNIST (FMNIST), SVHN, CIFAR-10, and CIFAR-100 dataset for 100 clients.
Dataset Overlap Dirichlet
$\alpha$
FL Alg. Second Session Transition Fourth Session Transition
Proposed Continuous Previous Average Proposed Continuous Previous Average
MNIST 0.0 0.3 MOON 92.04±0.24 76.92±0.34 57.82±1.23 64.19±0.85 93.08±0.17 81.93±0.42 69.35±0.75 77.92±0.72
FedProx 92.08±0.24 76.92±0.34 58.92±1.15 66.45±0.94 92.86±0.26 81.93±0.42 70.12±0.76 79.51±0.66
SCAFFOLD 90.69±0.42 84.30±3.26 50.36±1.91 49.97±2.04 91.95±0.39 86.39±2.63 64.93±1.35 68.55±1.45
FedACG 93.60±0.21 76.92±0.34 60.92±0.71 71.74±0.51 94.14±0.17 81.93±0.42 70.17±0.50 82.66±0.49
0.7 MOON 92.07±0.24 77.33±0.41 58.26±1.18 64.78±0.77 93.09±0.17 82.17±0.37 69.65±0.69 78.30±0.59
FedProx 92.10±0.24 77.33±0.41 59.20±1.13 66.76±0.89 92.88±0.26 82.17±0.37 70.33±0.73 79.72±0.58
SCAFFOLD 91.18±0.46 89.12±0.77 50.85±1.74 51.03±2.03 92.31±0.27 88.76±0.84 65.27±1.18 69.68±1.32
FedACG 93.60±0.21 77.33±0.41 61.15±0.65 72.02±0.50 94.16±0.18 82.17±0.37 70.36±0.49 82.83±0.46
0.2 0.3 MOON 92.58±0.16 80.98±0.27 70.49±0.97 72.59±0.97 93.56±0.17 84.65±0.21 79.21±0.63 82.68±0.54
FedProx 92.61±0.15 80.98±0.27 71.42±0.89 74.21±0.93 93.27±0.18 84.65±0.21 79.71±0.65 83.52±0.47
SCAFFOLD 91.32±0.30 86.65±3.84 64.11±1.24 61.30±2.11 92.50±0.30 87.27±1.50 75.88±0.75 75.74±0.95
FedACG 93.98±0.14 80.98±0.27 72.61±0.66 77.79±0.65 94.48±0.10 84.65±0.21 79.45±0.42 85.73±0.35
0.7 MOON 92.61±0.16 81.29±0.28 70.80±0.94 73.00±0.95 93.57±0.16 84.82±0.21 79.39±0.63 82.86±0.53
FedProx 92.63±0.14 81.29±0.28 71.66±0.88 74.42±0.92 93.28±0.19 84.82±0.21 79.82±0.65 83.66±0.47
SCAFFOLD 91.72±0.32 90.06±0.86 64.59±1.28 61.98±1.94 92.78±0.24 89.85±0.82 76.13±0.80 76.28±0.98
FedACG 93.99±0.15 81.29±0.28 72.81±0.68 77.97±0.68 94.51±0.13 84.82±0.21 79.57±0.43 85.83±0.36
FMNIST 0.0 0.3 MOON 83.13±0.09 59.36±0.90 77.09±0.40 77.84±0.49 84.88±0.19 64.58±0.25 79.48±0.29 80.04±0.42
FedProx 83.19±0.08 59.36±0.90 77.74±0.37 78.57±0.50 84.22±0.21 64.58±0.25 79.90±0.32 80.52±0.46
SCAFFOLD 77.80±0.42 65.94±5.97 47.74±0.88 47.93±1.88 79.88±0.17 74.69±4.63 53.64±0.54 60.72±0.93
FedACG 85.10±0.22 59.36±0.90 77.61±0.25 80.44±0.38 85.82±0.07 64.58±0.25 80.01±0.31 82.22±0.35
0.7 MOON 83.29±0.11 60.72±0.80 77.59±0.31 78.32±0.40 84.89±0.18 65.36±0.38 79.86±0.28 80.26±0.41
FedProx 83.33±0.13 60.72±0.80 78.16±0.35 78.93±0.43 84.29±0.15 65.36±0.38 80.28±0.34 80.72±0.38
SCAFFOLD 78.43±0.54 68.57±4.87 47.32±1.12 48.45±1.62 80.49±0.40 70.81±7.27 53.02±0.86 61.01±1.13
FedACG 85.17±0.25 60.72±0.80 78.03±0.29 80.68±0.39 85.81±0.13 65.36±0.38 80.45±0.26 82.41±0.29
0.2 0.3 MOON 82.88±0.15 72.53±0.29 74.83±0.31 76.78±0.59 84.30±0.14 75.78±0.16 77.15±0.17 79.48±0.53
FedProx 82.91±0.13 72.53±0.29 75.55±0.28 77.69±0.63 83.81±0.19 75.78±0.16 77.63±0.16 79.96±0.54
SCAFFOLD 77.35±0.53 69.90±6.23 55.82±1.59 50.39±4.48 79.45±0.21 70.79±4.93 61.12±1.40 61.10±2.25
FedACG 84.63±0.12 72.53±0.29 76.09±0.23 79.21±0.37 85.36±0.16 75.78±0.16 77.92±0.21 81.54±0.22
0.7 MOON 83.02±0.15 73.73±0.26 75.38±0.20 77.31±0.61 84.36±0.16 76.53±0.15 77.54±0.21 79.79±0.51
FedProx 83.04±0.15 73.73±0.26 75.99±0.19 78.14±0.73 83.95±0.22 76.53±0.15 78.06±0.21 80.22±0.56
SCAFFOLD 77.88±0.54 74.14±7.98 55.77±1.46 50.02±4.42 79.85±0.21 74.03±6.67 60.69±1.26 60.68±1.80
FedACG 84.70±0.14 73.73±0.26 76.45±0.22 79.65±0.37 85.41±0.13 76.53±0.15 78.31±0.16 81.82±0.25
SVHN 0.0 0.3 MOON 92.98±0.08 86.82±0.75 82.65±4.35 82.32±3.62 93.69±0.13 89.13±0.27 83.80±4.06 87.56±2.06
FedProx 91.78±0.16 86.54±0.18 86.87±2.00 88.26±0.36 92.36±0.12 87.93±0.28 89.25±1.10 90.59±0.38
SCAFFOLD 89.70±2.57 50.14±13.22 80.23±8.59 84.29±1.56 91.54±1.28 84.76±2.61 87.11±3.05 88.71±0.63
FedACG 93.09±0.16 86.56±0.23 80.67±2.41 86.95±2.77 93.48±0.27 87.95±0.28 75.82±10.48 90.30±1.67
0.7 MOON 93.21±0.18 88.72±0.33 84.88±3.19 87.03±3.18 93.95±0.13 90.37±0.21 87.69±1.98 90.44±1.50
FedProx 92.08±0.13 87.54±0.23 87.12±1.39 88.90±0.30 92.68±0.09 88.75±0.21 89.83±0.94 90.91±0.19
SCAFFOLD 92.06±0.61 68.61±11.01 86.32±3.01 85.88±1.23 93.13±0.45 90.40±1.04 90.28±1.10 89.47±0.86
FedACG 93.41±0.19 87.57±0.23 80.39±3.72 88.17±2.16 93.80±0.18 88.73±0.23 76.94±5.63 91.25±1.04
0.2 0.3 MOON 93.23±0.18 88.20±0.90 84.16±2.15 81.18±2.49 93.86±0.19 90.22±0.39 86.34±3.65 84.40±1.80
FedProx 92.10±0.21 87.44±0.81 88.04±2.09 85.45±1.25 92.75±0.21 89.05±0.50 90.01±1.32 88.57±0.57
SCAFFOLD 90.28±1.63 61.27±19.08 83.62±2.11 86.01±1.59 91.90±1.32 85.59±4.43 89.27±1.20 89.81±1.06
FedACG 93.34±0.25 87.43±0.84 84.39±1.47 89.51±1.49 93.66±0.20 89.06±0.48 87.97±1.96 91.29±0.80
0.7 MOON 93.54±0.24 89.92±0.32 85.92±2.37 85.59±2.71 94.24±0.19 91.20±0.22 89.17±1.73 88.30±1.71
FedProx 92.39±0.21 88.36±0.36 88.52±2.37 86.70±1.25 93.02±0.19 89.85±0.27 90.68±1.49 89.35±0.71
SCAFFOLD 92.10±0.82 75.97±8.03 87.49±1.29 86.91±0.86 93.30±0.56 89.57±1.58 90.79±0.46 90.60±0.20
FedACG 93.64±0.24 88.38±0.36 84.67±2.49 90.33±1.14 93.97±0.23 89.88±0.27 87.79±1.84 91.85±0.63
CIFAR-10 0.0 0.3 MOON 61.29±1.22 57.89±0.82 48.09±0.39 59.17±1.51 64.73±0.71 63.34±0.53 53.91±0.92 61.77±1.53
FedProx 61.40±1.09 57.96±0.68 45.94±0.76 59.18±1.37 64.06±0.75 63.38±0.57 51.16±0.56 61.91±1.31
SCAFFOLD 68.44±0.73 68.93±1.91 53.67±0.64 65.15±1.10 70.49±0.59 72.52±1.55 55.75±0.87 67.62±0.95
FedACG 61.88±2.33 57.88±0.75 44.12±1.29 64.33±1.53 64.05±2.01 63.33±0.40 45.60±1.78 65.57±1.50
0.7 MOON 66.28±0.51 61.39±0.73 49.03±1.55 63.70±0.46 69.46±0.45 67.11±0.51 57.95±1.08 66.00±0.53
FedProx 66.30±0.64 61.54±0.71 46.50±1.17 64.01±0.67 69.13±0.78 67.09±0.44 54.31±1.17 66.56±0.54
SCAFFOLD 72.95±0.68 71.91±0.81 57.37±0.84 69.29±0.86 75.12±0.81 76.43±0.38 59.80±0.57 71.71±0.89
FedACG 70.35±0.80 61.46±0.74 47.91±1.35 68.17±0.50 73.23±0.60 67.09±0.41 49.10±1.29 69.73±0.33
0.2 0.3 MOON 54.39±1.21 51.64±0.93 51.54±1.02 53.65±0.92 59.17±1.25 57.26±0.93 56.34±0.66 56.85±0.84
FedProx 55.26±1.80 51.74±0.88 50.40±0.99 53.29±0.80 59.97±1.44 57.42±0.74 54.30±0.96 56.91±0.68
SCAFFOLD 67.35±0.65 63.67±3.23 52.91±1.16 62.49±0.88 69.36±0.73 68.02±2.04 54.60±1.12 65.57±0.90
FedACG 59.33±2.61 51.48±0.88 48.80±1.07 55.33±0.99 62.01±2.09 57.31±0.67 49.07±2.10 59.35±0.73
0.7 MOON 61.78±1.59 57.64±0.52 53.14±1.17 58.49±0.82 66.10±1.53 64.19±0.45 59.72±0.74 60.73±0.73
FedProx 63.81±1.31 57.73±0.50 52.35±0.95 58.70±0.71 66.36±1.13 64.24±0.44 58.05±0.72 61.36±0.66
SCAFFOLD 71.33±0.81 69.56±0.60 58.65±0.59 67.46±0.45 73.51±0.74 74.41±1.15 61.05±0.85 70.39±0.38
FedACG 66.27±1.49 57.67±0.56 51.04±0.52 61.61±0.39 69.04±1.37 64.20±0.48 53.47±0.73 64.66±0.37
CIFAR-100 0.0 0.3 MOON 38.18±0.43 32.07±0.42 38.10±0.58 34.75±0.45 41.89±0.53 38.62±0.50 44.49±0.63 38.61±0.38
FedProx 37.79±0.37 31.95±0.36 37.05±0.85 34.16±0.40 41.55±0.34 38.65±0.57 43.55±0.94 38.08±0.57
SCAFFOLD 44.48±0.43 45.09±0.96 43.80±0.69 40.70±0.53 48.08±0.76 50.18±0.48 49.73±0.32 44.11±0.37
FedACG 46.66±0.28 31.96±0.28 31.38±0.77 40.49±0.87 49.69±0.47 38.68±0.38 42.63±0.67 44.33±0.77
0.7 MOON 40.64±0.70 35.09±0.46 39.90±0.94 36.73±0.63 44.58±0.96 41.55±0.49 45.92±1.56 40.46±0.59
FedProx 40.42±0.46 35.15±0.51 38.29±1.45 36.47±0.50 44.01±0.60 41.60±0.50 44.66±0.96 40.57±0.54
SCAFFOLD 47.18±0.50 46.95±0.59 46.47±0.53 42.65±0.49 50.34±0.57 50.85±0.59 52.22±0.85 46.53±0.37
FedACG 48.81±0.57 35.09±0.50 32.81±0.62 42.18±0.83 51.58±0.90 41.74±0.38 43.66±1.19 45.83±0.51
0.2 0.3 MOON 44.29±0.74 39.28±0.87 44.60±0.72 36.83±0.81 47.17±0.98 45.51±0.33 50.41±0.81 40.36±0.45
FedProx 43.85±0.92 39.65±0.45 45.31±0.84 38.71±0.31 47.01±0.78 45.51±0.28 50.48±0.46 42.39±0.38
SCAFFOLD 49.14±0.41 48.24±1.03 47.23±1.12 42.93±0.41 52.06±0.35 51.13±0.65 52.32±1.47 46.33±0.37
FedACG 51.17±0.25 39.52±0.81 43.37±1.28 43.78±1.05 53.16±0.31 45.04±0.71 49.33±0.89 48.11±1.20
0.7 MOON 46.86±0.45 42.32±0.50 46.13±0.30 39.05±0.76 50.18±0.35 47.98±0.52 52.30±0.48 42.75±0.61
FedProx 46.31±0.45 42.50±0.52 47.57±0.22 40.55±0.40 49.34±0.48 48.09±0.42 52.63±0.32 44.13±0.50
SCAFFOLD 50.91±0.79 50.95±0.71 50.99±0.44 46.96±0.45 53.82±1.07 54.19±0.50 56.22±0.20 50.67±0.40
FedACG 52.97±0.39 42.32±0.55 44.94±0.92 47.48±0.66 55.20±0.45 48.11±0.30 51.74±0.71 51.15±0.47
Table 6 further illustrates the effectiveness of Algorithm 1 on TinyImagenet for a system of 100 clients.
Table 6: Performance comparison of implementing Algorithm 1 with MOON, FedProx, SCAFFOLD and FedACG under different label distributions for TinyImagenet dataset for a system with 100 clients.
Dataset Overlap Dirichlet
$\alpha$
FL Alg. First Session Transition Second Session Transition
Proposed Continuous Previous Average Proposed Continuous Previous Average
TinyImagenet 0.0 0.3 MOON 65.98±0.30 56.16±0.26 58.58±0.51 58.15±0.33 67.03±0.42 57.64±0.67 61.71±0.79 62.57±0.45
FedProx 65.33±0.30 56.03±0.31 63.03±0.36 61.15±0.38 66.56±0.41 57.57±0.48 64.73±0.14 63.80±0.22
SCAFFOLD 67.51±0.17 52.43±0.92 60.19±0.26 61.66±0.24 68.47±0.23 54.64±0.65 61.11±0.38 63.81±0.27
FedACG 66.54±0.19 56.05±0.34 32.99±3.50 55.69±1.47 67.11±0.47 57.78±0.50 36.26±1.55 62.29±0.35
0.7 MOON 66.29±0.47 57.73±0.29 59.72±0.70 61.17±0.57 67.77±0.32 58.79±0.42 62.64±0.55 64.33±0.33
FedProx 66.21±0.21 57.71±0.35 63.88±0.22 61.98±0.37 67.26±0.16 58.75±0.46 65.25±0.39 64.40±0.21
SCAFFOLD 68.55±0.35 55.33±0.63 60.67±0.54 62.50±0.32 69.23±0.41 56.46±0.41 61.12±0.85 64.71±0.28
FedACG 66.71±0.43 57.74±0.29 38.89±2.21 58.76±1.20 67.84±0.28 58.78±0.41 31.27±21.23 63.96±0.30
0.2 0.3 MOON 64.99±0.66 58.67±0.35 60.88±0.51 60.79±0.29 66.36±0.61 59.79±0.42 61.91±0.54 62.98±0.27
FedProx 65.48±0.43 58.64±0.26 63.75±0.17 60.47±0.35 66.32±0.32 59.80±0.43 65.38±0.28 63.45±0.50
SCAFFOLD 67.58±0.46 56.77±0.49 60.84±0.45 62.22±0.38 68.09±0.31 57.66±0.69 61.35±0.57 64.48±0.45
FedACG 66.33±0.70 58.65±0.35 57.24±0.51 58.62±0.42 66.39±0.40 59.84±0.48 55.29±0.96 62.42±0.49
0.7 MOON 66.09±0.58 59.94±0.41 62.47±0.20 62.56±0.28 67.03±0.39 60.69±0.52 63.18±0.49 64.20±0.22
FedProx 66.06±0.22 60.00±0.51 64.40±0.17 61.49±0.31 67.06±0.31 60.58±0.54 65.46±0.22 64.06±0.38
SCAFFOLD 68.30±0.19 58.53±0.50 61.07±0.41 63.57±0.25 68.83±0.33 58.76±0.41 61.16±0.28 65.35±0.40
FedACG 66.51±0.34 60.02±0.42 56.94±0.73 61.96±0.45 67.06±0.12 60.58±0.44 56.75±0.72 63.95±0.36
Table 7 further illustrates the effectiveness of Algorithm 1 on AG News dataset for a system of 100 clients.
Table 7: Performance comparison of implementing Algorithm 1 with MOON, FedProx, SCAFFOLD and FedACG under different label distributions for AG News dataset for 100 clients.
Dataset Overlap Dirichlet
$\alpha$
FL Alg. First Session Transition Second Session Transition
Proposed Previous Average Proposed Previous Average
AG News 0.0 0.3 MOON 73.94±1.43 14.04±0.29 14.22±0.23 75.19±1.35 14.34±0.31 24.60±2.00
FedProx 73.89±1.43 14.03±0.28 14.20±0.21 75.10±1.41 14.34±0.31 24.33±1.95
SCAFFOLD 66.46±2.18 12.66±0.36 12.86±0.18 69.33±1.78 13.44±0.45 13.85±0.25
FedACG 76.27±1.28 14.52±0.30 16.95±1.06 77.18±1.41 14.78±0.29 35.19±2.65
0.7 MOON 73.95±1.40 14.04±0.30 14.23±0.23 74.83±1.50 14.34±0.32 24.59±1.99
FedProx 73.90±1.40 14.03±0.29 14.20±0.21 74.75±1.56 14.34±0.31 24.32±1.93
SCAFFOLD 66.64±2.14 12.63±0.21 12.88±0.18 68.99±3.04 13.24±0.16 13.89±0.26
FedACG 76.25±1.28 14.52±0.29 16.96±1.06 76.88±1.55 14.78±0.29 35.19±2.64
0.8 0.3 MOON 59.15±0.66 51.98±0.85 53.36±0.85 60.78±1.19 52.97±0.83 53.94±0.97
FedProx 59.05±0.62 51.96±0.83 53.34±0.85 60.54±1.31 52.96±0.83 53.93±0.96
SCAFFOLD 47.26±0.62 48.35±0.39 48.58±1.07 49.11±1.25 50.09±0.52 49.48±0.96
FedACG 63.95±0.28 52.89±0.76 55.26±0.66 65.28±0.90 53.83±0.57 55.82±0.51
0.7 MOON 59.20±0.67 51.99±0.83 53.36±0.85 60.48±0.26 52.97±0.83 53.95±0.98
FedProx 59.09±0.65 51.97±0.83 53.34±0.84 60.27±0.34 52.96±0.83 53.92±0.98
SCAFFOLD 47.37±0.61 48.87±0.21 48.73±1.07 48.68±0.72 50.42±0.27 49.58±1.00
FedACG 63.95±0.29 52.92±0.75 55.26±0.66 64.54±0.36 53.84±0.59 55.83±0.50
A.2 Additional Results for Alternative Label Distributions
In addition to the label distribution we used in the main text, we also evaluate Algorithm 1 using the following distribution.
  • Two-Shard [11, 48, 12, 49, 13]: For datasets with 10 labels, each label is divided into two equally-sized shards, and each client receives two different label shards. For datasets with 100 labels, the labels are divided into 10 non-overlapping batches, and each batch is split into two shards, which are then assigned to two randomly selected clients. Each client ends up with data from two labels for 10-label datasets, or 20 labels for 100-label datasets .
  • Half: Half the clients have one half of the labels, and the rest have the other half. Each client has all the labels from their assigned half (i.e., 5 labels for 10-label datasets or 50 labels for 100-label datasets), and the data is evenly distributed, meaning that the amount of data for each label is equally divided among the clients.
  • Partial-Overlap: Two sets of labels are selected, each containing 60% of the total labels, with a 20% overlap between them. Each client in the first set has 6 labels for 10-label datasets (or 60 labels for 100-label datasets), and each client in the second set has a similar distribution. The overlapping labels are split between the two halves of clients, with half of the data for overlapping labels going to the first client set and the other half to the second set. The non-overlapping labels are assigned to the clients within each set, and the data corresponding to these labels is evenly distributed across the clients, meaning that each client receives an approximately equal share of the data for their assigned labels.
  • Distinct: Each client is assigned a unique set of labels. For datasets with 10 labels, each client receives 1 unique label, while for datasets with 100 labels, each client receives 10 unique labels.
Figure 3 highlights selected scenarios and illustrates the advantages of the proposed algorithm for all FL algorithms during periods of pronounced performance drops or gains due to data distribution shifts. The algorithm prioritizes models trained on past distributions similar to the current one, enabling the initial model to adapt more quickly and consistently outperform baseline approaches. These results underscore the robustness and effectiveness of the proposed approach across diverse FL algorithms, datasets, models, and dynamic data distributions.
Figure 3: Performance comparison of the proposed algorithm implemented with four FL algorithms under “Two-Shard” and “Partial-Overlap” label distributions across selected datasets. The results demonstrate the robustness of our proposed scheme to dynamic data distributions caused by client arrivals and departures.
Table 8 compares the average accuracy of the proposed algorithm across all FL algorithms during the first 10 global rounds after three data distribution shifts under the “Two-Shard” and “Partial-Overlap” label distributions. The results across datasets, label distributions, and models show that the proposed algorithm mitigates performance degradation from dynamic client arrivals and departures. Moreover, it accelerates performance recovery when the current data distribution aligns with a previously observed one.
Table 8: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and MOON under different label distributions and datasets. Performance is measured across 3 transitions for each dataset.
FL Algorithm Label Distribution Dataset (Model) 1st Transition 2nd Transition 3rd Transition
Proposed Previous Average Proposed Previous Average Proposed Previous Average
FedProx Two-Shard MNIST (MLP) 98.45 96.4 98.24 94.77 90.72 93.19 98.56 96.53 98.4
Fashion-MNIST (MLP) 92.04 90.31 88.89 97.59 96.38 97.33 95.39 93.54 94.58
SVHN (CNN) 86.52 40.13 72.7 97.44 84.53 95.08 95.8 55.16 83.08
CIFAR10 (ResNet18) 41.64 35.32 39.68 46.26 35.62 44.62 82.58 58.34 78.3
CIFAR100 (ResNet18) 58.46 56.68 51.07 49.28 46.52 47.7 52.73 48.35 51.49
Partial-Overlap MNIST (MLP) 62.88 56.16 56.23 64.27 57.14 53.96 63.53 57.41 56.81
Fashion-MNIST (MLP) 87.06 83.79 86.78 88.06 85.52 84.78 87.2 84.33 86.89
SVHN (CNN) 92.4 76.09 91.72 93.64 68.67 92.35 92.51 71.79 92.02
CIFAR10 (ResNet18) 81.5 77.05 77.31 86.99 83.88 84.97 81.69 79.84 77.62
CIFAR100 (ResNet18) 59.09 54.94 52.23 59.22 55.09 56.13 59.34 56.11 53.34
FedAvg Two-Shard MNIST (MLP) 98.45 96.4 98.24 94.77 90.72 93.19 98.56 96.53 98.4
Fashion-MNIST (MLP) 92.04 90.31 88.89 97.59 96.38 97.33 95.39 93.54 94.58
SVHN (CNN) 76.53 51.03 56.65 83.45 39.58 72.65 95.87 42.97 85.15
CIFAR10 (ResNet18) 41.98 32.77 41.05 51.01 36.57 48.57 82.51 64.03 81.04
CIFAR100 (ResNet18) 47.88 44.29 46.87 49.63 46.04 47.64 52.18 48.43 51.17
Partial-Overlap MNIST (MLP) 93.81 89.65 89.33 91.11 87.46 90.35 94.16 90.1 90.67
Fashion-MNIST (MLP) 87.58 85.06 83.04 86.77 83.88 86.49 87.64 85.59 84.16
SVHN (CNN) 92.27 79.11 89.94 93.67 82.01 91.46 92.48 77.2 91.47
CIFAR10 (ResNet18) 85.87 82.35 83.84 79.94 77.1 75.31 85.87 82.9 84.41
CIFAR100 (ResNet18) 56.94 53.29 54.31 57.25 54.17 51.71 57.35 54.41 54.6
SCAFFOLD Two-Shard MNIST (MLP) 97.97 51.69 96.73 92.27 42.16 44.86 98.32 54.78 97.26
Fashion-MNIST (MLP) 84.24 36.59 22.94 96.89 51.82 96.18 92.29 52.41 45.45
SVHN (CNN) 79.23 69.2 47.94 95.32 90.76 73.11 93.96 88.37 73.25
CIFAR10 (ResNet18) 46.63 10.74 23.31 92.39 38.32 90.59 56.11 12.66 24.98
CIFAR100 (ResNet18) 21.63 5.61 9.59 35.75 21.3 19.42 26.18 7.69 11.98
Partial-Overlap MNIST (MLP) 85.72 61.6 71.39 89.66 77.03 53.64 87.1 69.45 72.93
Fashion-MNIST (MLP) 86.67 62.41 74.68 87.54 63.48 68.97 87.64 65.2 74.91
SVHN (CNN) 90.33 84.02 73.88 91.9 89.05 80.91 91.33 85.83 74.68
CIFAR10 (ResNet18) 72.64 37.95 42.06 79.12 39.31 69.88 74.52 40.29 49.57
CIFAR100 (ResNet18) 61.03 52.56 48.94 61.16 53.54 58.66 60.79 53.99 53.2
MOON Two-Shard MNIST (MLP) 98.15 96.54 98.04 98.38 96.55 98.24 98.52 96.48 98.39
Fashion-MNIST (MLP) 92.08 90.28 88.84 97.42 96.47 97.31 95.36 93.54 94.53
SVHN (CNN) 87.0 78.61 75.72 69.51 64.57 54.88 93.49 90.2 89.87
CIFAR10 (ResNet18) 44.03 40.53 39.42 51.31 43.46 41.82 85.0 71.45 73.54
CIFAR100 (ResNet18) 57.94 56.05 44.18 45.8 43.81 44.71 62.06 60.56 53.54
Partial-Overlap MNIST (MLP) 93.85 89.65 89.33 91.1 87.64 90.35 94.12 90.2 90.67
Fashion-MNIST (MLP) 87.6 85.06 83.04 86.66 83.66 86.49 88.01 85.53 84.16
SVHN (CNN) 92.96 90.41 88.39 91.74 87.97 90.77 93.18 91.25 90.25
CIFAR10 (ResNet18) 85.53 83.59 84.67 79.28 77.57 76.04 79.37 78.44 76.11
CIFAR100 (ResNet18) 55.52 52.58 47.77 55.27 52.22 52.7 55.72 53.49 49.01
Table 9: Performance comparison of FedProx, FedAvg, SCAFFOLD, and MOON under “Half” and “Distinct” label distributions across various datasets. Performance is evaluated over three transitions for each dataset.
FL Algorithm Label Distribution Dataset (Model) 1st Transition 2nd Transition 3rd Transition
Proposed Previous Average Proposed Previous Average Proposed Previous Average
FedProx Half MNIST (MLP) 96.02 90.12 92.45 96.29 90.56 93.59 93.25 88.17 92.45
Fashion-MNIST (MLP) 90.97 86.02 90.17 86.62 85.06 84.2 91.29 86.67 90.26
SVHN (CNN) 93.23 27.0 92.18 91.64 17.15 90.98 93.53 69.5 92.75
CIFAR10 (ResNet18) 86.32 48.83 83.59 90.38 54.74 87.64 87.68 58.41 84.62
CIFAR100 (ResNet18) 62.88 56.16 56.23 64.27 57.14 53.96 63.53 57.41 56.81
Distinct MNIST (MLP) 98.05 92.75 97.71 95.98 88.77 93.96 94.11 85.66 90.54
Fashion-MNIST (MLP) 95.5 92.4 95.0 95.8 90.59 94.56 94.02 87.36 90.73
SVHN (CNN) 92.25 45.18 67.51 87.91 38.45 63.85 92.24 42.7 60.68
FedAvg Half MNIST (MLP) 96.03 90.12 92.45 92.56 87.83 92.45 96.21 90.56 93.59
Fashion-MNIST (MLP) 86.24 84.54 82.54 90.29 86.12 89.91 90.5 86.62 90.1
SVHN (CNN) 91.63 71.49 88.78 93.54 51.98 91.07 91.98 42.34 90.55
CIFAR10 (ResNet18) 90.19 70.15 88.08 86.23 77.41 81.67 90.03 73.55 89.25
CIFAR100 (ResNet18) 57.41 52.95 54.37 57.09 53.7 51.07 57.48 53.38 54.44
Distinct MNIST (MLP) 97.97 93.03 97.57 95.72 89.19 93.65 94.04 86.11 90.4
Fashion-MNIST (MLP) 95.81 90.66 94.44 93.88 87.5 90.56 95.8 92.0 94.86
SVHN (CNN) 94.99 49.88 80.62 91.9 38.94 58.53 87.21 32.71 66.47
SCAFFOLD Half MNIST (MLP) 85.53 29.8 82.93 92.58 57.81 30.18 89.08 51.43 84.39
Fashion-MNIST (MLP) 77.18 36.71 28.55 83.35 65.31 77.91 79.51 40.82 42.19
SVHN (CNN) 91.66 86.87 76.06 89.86 84.79 80.25 93.04 87.31 87.18
CIFAR10 (ResNet18) 78.63 17.35 34.76 84.62 19.59 79.32 80.13 20.86 49.7
CIFAR100 (ResNet18) 28.92 9.97 18.74 31.11 13.3 6.44 34.23 13.84 21.62
Distinct MNIST (MLP) 96.95 50.5 95.35 93.08 22.24 60.96 91.36 21.06 22.78
Fashion-MNIST (MLP) 94.08 32.05 91.66 93.28 13.81 53.5 91.72 29.1 33.07
MOON Half MNIST (MLP) 96.06 90.16 92.45 92.64 87.78 92.45 96.29 90.46 93.59
Fashion-MNIST (MLP) 86.12 84.43 82.54 90.33 85.79 89.91 86.31 84.96 83.75
SVHN (CNN) 93.64 89.35 89.95 92.11 87.3 91.18 93.63 90.54 91.37
CIFAR10 (ResNet18) 89.0 85.77 88.26 84.89 81.08 81.53 89.16 86.67 88.51
CIFAR100 (ResNet18) 54.9 52.27 53.8 55.95 52.75 48.99 55.05 52.63 54.07
Distinct MNIST (MLP) 95.67 88.67 93.56 94.07 85.7 90.2 95.63 90.65 94.32
Fashion-MNIST (MLP) 95.76 90.14 94.4 93.86 86.83 90.06 95.66 91.76 94.85
Table 9 presents a comparative analysis of the average accuracy achieved by the proposed algorithm across all federated learning (FL) algorithms during the first 10 global rounds following three shifts in data distribution under the “Half” and “Distinct” label distributions.
Table 10 presents the results of applying Algorithm 1 with FedAvg for 20 clients on the TinyImageNet dataset [50]. The baseline method used is “Average.” These results highlight the effectiveness of Algorithm 1 in handling scenarios with a larger client population and a more complex dataset using a more advanced model ResNet34.
Table 10: Performance of Algorithm 1 with FedAvg under “Half” and “Partial-Overlap” label distributions for the TinyImageNet dataset with 20 clients. Performance is measured across three transitions for each dataset.
Label Distribution Dataset (Model) 1st Transition 2nd Transition 3rd Transition
Proposed Previous Proposed Previous Proposed Previous
Half TinyImageNet (ResNet34) 80.72 72.81 77.77 72.24 80.43 73.42
Partial-Overlap TinyImageNet (ResNet34) 92.27 79.11 93.67 82.01 92.48 77.2
Table 11 shows the results for 100 clients for Algorithm 1 integrated with FedProx, FedAvg, SCAFFOLD, and MOON on the MNIST and Fashion-MNIST datasets, using 100 clients under the Partial-Overlap label distribution. Performance is evaluated over three transitions for each dataset.
Table 11: Performance comparison of Algorithm 1 integrated with FedProx, FedAvg, SCAFFOLD, and MOON on the MNIST and Fashion-MNIST datasets, utilizing 100 clients for FedProx, FedAvg, and MOON, and 50 clients for SCAFFOLD under the Partial-Overlap label distribution. Performance is evaluated across three transitions for each dataset.
FL Algorithm Dataset (Model) 1st Transition 2nd Transition 3rd Transition
Proposed Previous Average Proposed Previous Average Proposed Previous Average
FedProx MNIST (MLP) 94.09 89.53 89.35 91.36 87.4 90.45 94.33 89.92 90.78
Fashion-MNIST (MLP) 87.88 85.13 83.37 87.14 83.86 86.76 88.08 85.57 84.65
FedAvg MNIST (MLP) 90.86 86.15 90.25 94.09 89.53 89.35 91.36 87.4 90.45
Fashion-MNIST (MLP) 86.55 82.78 86.4 87.88 85.13 83.37 87.14 83.86 86.76
SCAFFOLD MNIST (MLP) 64.82 33.15 52.67 82.67 62.47 30.18 74.04 50.77 59.66
MOON MNIST (MLP) 91.36 87.4 90.45 94.33 89.92 90.78 91.55 87.87 90.55
Fashion-MNIST (MLP) 87.14 83.86 86.75 88.07 85.57 84.65 87.31 84.34 86.9
Figure 4 highlights the performance of Algorithm 1 in more selected scenarios and illustrates the advantages of the proposed algorithm for all FL algorithms during periods of pronounced performance drops or gains due to data distribution shifts. The algorithm prioritizes models trained on past distributions that share similarities with the current one, enabling the initial model to adapt more quickly to the new distribution and consistently outperforming baseline approaches. These results underscore the robustness and effectiveness of the proposed approach across diverse FL algorithms, datasets, models, and dynamic data distributions.
Figure 4: Performance comparison of the proposed algorithm implemented with four FL algorithms under all label distributions across more selected datasets. The results demonstrate the robustness of our proposed scheme to dynamic data distributions caused by client arrivals and departures.
Appendix B Gradient Similarity, Sensitivity Analysis and Ablation Studies for Algorithm 1
B.1 Computed Gradient and Similarity
In our method, the computed gradient is the normalized global gradient rather than any single client’s raw gradient. During aggregation, each client’s local gradient is weighted by the size of its local dataset, ensuring that clients with more data contribute proportionally more. Consequently, the global gradient captures collective data heterogeneity without being biased by client-size.
The utility of gradient similarity for model initialization depends on the alignment between current and historical data distributions. When client distributions evolve gradually and exhibit recurring patterns—a typical scenario in federated learning—gradient similarity provides an effective heuristic for selecting an initial model that adapts rapidly to new clients. While our method excels under moderate to high overlap conditions, completely novel classes still pose an inherent challenge in non-stationary environments.
To assess robustness under extreme shifts, we introduced unseen labels only in the final session and measured average test accuracy over the first ten global rounds. As shown in Table 12, our initialization strategy maintains performance comparable to baseline approaches.
Table 12: Average accuracy over the first ten rounds under the unseen-label edge case.
Algorithm Dataset Proposed Previous Average
FedAvg Fashion-MNIST 94.98 94.75 95.89
MOON SVHN 94.17 89.53 94.02
B.2 Sensitivity Analysis of the Similarity Scaling Factor $R$
We evaluate the effect of the scaling factor $R$ in the initial-model construction (5) on both FedAvg and FedProx. Accuracy across multiple transitions remains stable for a wide range of $R$ values.
Table 13: Sensitivity of transition performance to the similarity-scaling factor $R$ for FedAvg and FedProx on SVHN and CIFAR-100.
Algorithm Dataset $\mathbf{R}$ 1st Transition 2nd Transition 3rd Transition
FedAvg SVHN 50 91.78 93.49 91.87
SVHN 80 91.76 93.58 91.90
SVHN 100 91.63 93.54 91.98
SVHN 120 91.80 93.59 91.83
SVHN 140 91.77 93.47 91.75
CIFAR-100 110 58.36 56.63 58.37
CIFAR-100 120 56.95 56.28 57.06
CIFAR-100 130 57.41 56.29 57.46
CIFAR-100 140 58.30 56.37 58.42
CIFAR-100 150 57.46 56.75 58.05
FedProx SVHN 50 92.43 93.83 92.54
SVHN 80 92.44 93.88 92.62
SVHN 100 92.40 93.64 92.51
SVHN 120 92.49 93.79 92.71
SVHN 140 92.50 93.71 92.70
CIFAR-100 110 58.55 58.70 58.57
CIFAR-100 120 58.56 58.63 58.78
CIFAR-100 130 59.02 58.50 59.06
CIFAR-100 140 58.93 58.88 59.07
CIFAR-100 150 59.66 59.51 59.68
B.3 Ablation Studies
We have conducted additional experiments replacing the pilot model with a randomly initialized one to compute gradients. Performance drops in most cases, supporting the utility of pilot model knowledge.
Table 14: Ablation study comparing the proposed pilot model versus a randomly initialized pilot across transitions and settings.
Setting 1st Transition 2nd Transition 3rd Transition
Proposed Random Pilot Proposed Random Pilot Proposed Random Pilot
MOON, Two-shards, SVHN 87.00 86.03 69.51 63.14 93.49 93.32
SCAFFOLD, HALF, SVHN 91.66 71.78 89.86 87.54 93.04 84.67
FedProx, Slight Overlap, CIFAR10 81.50 79.47 86.99 86.80 81.69 79.62
MOON, Slight Overlap, SVHN 92.96 91.91 91.74 91.60 93.18 91.74
Appendix C Computation, Communication, and Storage Analysis of Algorithm 1
C.1 Cost Analysis with Variable Clients
The breakdown below provides a precise view of costs per phase. Notation follows the main text; session indexing will be refined in the revised manuscript.
Pilot Model Training (consisting of $P$ sessions):
  • Computation: Each of the $P$ pilot sessions comprises $T$ global rounds. In round $g$ ($1\leq g\leq T$) of a pilot session, each client $k\in\mathcal{K}^{(g)}$ executes $e_{k}^{(g)}$ local epochs. Total local epochs:
    $P\,\times\,\sum_{g=1}^{T}\sum_{k\in\mathcal{K}^{(g)}}e_{k}^{(g)}.$
  • Communication: There are $P\times T$ global rounds. In each round $g$, $|\mathcal{K}^{(g)}|$ clients upload updates and the server broadcasts the aggregated model. Total transmissions:
[…]
  • Storage: Model size is $W$. We temporarily store $P$ intermediate models of size $W$, which are discarded after constructing the pilot model.
Gradient Computation (for session $s$, $s\in\{P,\dots,S\}$, at global round $g=sT$):
  • Computation: Perform $V$ additional global rounds initialized from the pilot model. Clients in $\mathcal{K}^{(sT)}$ each run $e_{k}^{(sT)}$ local epochs per round, for a total of
    $V\,\times\,\sum_{k\in\mathcal{K}^{(sT)}}e_{k}^{(sT)}.$
  • Communication: Over $V$ rounds, each round incurs $|\mathcal{K}^{(sT)}|$ uploads and one broadcast. Total:
[…]
  • Storage: One gradient of size $W$ is stored for session $s$.
Coefficient Calculation (for session $m$, $m\in\{P+1,\dots,S\}$, at $g=mT$):
  • Computation: Computing softmax-based coefficients over the $(m-P)$ stored items incurs
    $\mathcal{O}(m-P)$
    cost for vector norms and exponentials.
Total Persistent Storage:
  • Pilot model: One model of size $W$ (intermediate checkpoints discarded).
  • Per post-pilot session $s$:
    • Final model checkpoint: $W$,
    • Computed gradient: $W$.
  • Overall growth: Linear increase of $2W$ per post-pilot session, scaling with $(S-P)$.
C.2 Storage Overhead
Persistent storage grows linearly with $S-P$ (one model and one gradient per session), which remains practical.
C.3 Justification for Computation and Communication Requirements
The upfront cost of pilot training and gradient computation is offset by accelerated adaptation to shifting client distributions, reducing the total number of global rounds (e.g., avoiding $50+$ recovery rounds required by standard baselines).
C.4 Mitigating Storage for Matrices $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$
To bound storage growth, one may:
  • Limited history: Retain only the most recent $H$ sessions’ models/gradients.
  • Selective retention: Store models/gradients that maximize distributional diversity, preserving initialization quality with fewer checkpoints.
These strategies ensure manageable storage while maintaining algorithmic efficacy in dynamic federated learning settings.
Appendix D Experimental Setup, Reproducibility, and Computational Resources
D.1 Setup and Reproducibility
The following command was used to launch each experiment. The meaning and purpose of each argument are described in detail in our accompanying README.md.
python -u fasy_adapt_training.py \
    --dataset_name ’cifar10’ \
    --algorithm ’scaffold’ \
    --cross_session_label_overlap 0.0 \
    --in_session_label_dist "dirichlet" \
    --dirichlet_alpha 0.7 \
    --seed 300 \
    --initial 1 \
    --average 1 \
    --previous 1 \
    --continuous 0 \
    --num_clients 100 \
    --num_sessions 7 \
    --num_sessions_pilot 1 \
    --num_rounds_pilot 100 \
    --num_rounds_actual 100 \
    --lr_config_path ’lr_poly.json’ \
    --momentum 0.9 \
    --gpu_index 0 \
    --num_SGD_training 5 \
    --batch_size_training 128 \
    --num_SGD_grad_cal 5 \
    --batch_size_grad_cal 128 \
    --prox_alpha 1.0 \
    --moon_mu 1.0 \
    --moon_tau 1.0 \
    --kl_coefficient 0.5 \
    --acg_beta 0.1 \
    --acg_lambda 0.5 \
    --num_round_grad_cal 1 \
    --similarity "two_norm" \
    --similarity_scale 10.0
How hyperparameters were chosen: We selected all hyperparameters such that the training within each session converges by its final round. This ensures a fair comparison across different algorithms and datasets.
Baseline variants: AG News uses the average and previous baselines. All other datasets use average, previous, and continuous baselines.
20 Clients Configuration
  • num_sessions=6
  • acg_beta=0.01, acg_lambda=0.85
  • prox_alpha=1.0 for FedProx
  • Algorithms:
    • All datasets: FedAvg, FedProx, SCAFFOLD, FedACG
    • TinyImageNet: MOON (moon_mu=1.0, moon_tau=1.0), FedProx, FedACG
  • kl_coefficient = 1.0 for all datasets except AG News (0.5)
  • similarity_scale = 10.0 for all datasets except AG News (800)
  • Random seeds: 0, 1, 2 (TinyImageNet uses 100, 200, 300)
100 Clients Configuration
  • num_sessions=7
  • acg_beta=0.1, acg_lambda=0.5
  • Algorithms:
    • All datasets: MOON (moon_mu=1.0, moon_tau=1.0), FedProx (prox_alpha=1.0), SCAFFOLD, FedACG
  • kl_coefficient = 1.0 for all datasets except AG News (0.5)
  • similarity_scale = 10.0 for image datasets; AG News uses 800
  • Random seeds: 100, 200, 300
Training Rounds (all clients)
  • MNIST, Fashion-MNIST: 50 pilot rounds, 50 actual rounds
  • SVHN, TinyImageNet: 200 pilot rounds, 200 actual rounds
  • CIFAR-10, CIFAR-100: 150 pilot rounds, 150 actual rounds
  • AG News: 100 pilot rounds, 100 actual rounds
Shared Parameters Across All Experiments
  • momentum=0.9
  • num_SGD_training=5, batch_size_training=128
  • num_SGD_grad_cal=5, batch_size_grad_cal=128
  • num_round_grad_cal=1
  • Learning rate scheduler: lr_config_path = ’lr_poly.json’
D.2 Compute Resources
All experiments were conducted on a high-performance node provided by our university’s Advanced Computing Center, equipped with two Intel Xeon Platinum 8480+ 56-core CPUs (3.8 GHz), eight NVIDIA H100 GPUs (80 GB each), and 1031 GB of RAM. Each SLURM job was allocated one H100 GPU and 14 CPU cores.
LABEL:table:resource_20_appendix presents a comprehensive overview of average CPU time and maximum memory usage for our experiments with 20 clients. The average is taken over three random seeds.
Table 15: Average CPU time and average maximum memory usage across three random seeds for our experiments with 20 clients.
Dataset Cross Alpha Method CPU Time Max Mem
MNIST $0.2$ $0.3$ FedAvg 00:26:28 1.07G
MNIST $0.2$ $0.3$ FedProx 00:26:33 1.11G
MNIST $0.2$ $0.3$ SCAFFOLD 00:26:19 1.06G
MNIST $0.2$ $0.3$ FedACG 00:26:18 1.10G
MNIST $0.2$ $0.7$ FedAvg 00:25:51 1.06G
MNIST $0.2$ $0.7$ FedProx 00:26:00 1.11G
MNIST $0.2$ $0.7$ SCAFFOLD 00:25:45 1.05G
MNIST $0.2$ $0.7$ FedACG 00:26:09 1.11G
MNIST $0.0$ $0.3$ FedAvg 00:26:04 1.05G
MNIST $0.0$ $0.3$ FedProx 00:26:13 1.12G
MNIST $0.0$ $0.3$ SCAFFOLD 00:25:49 1.06G
MNIST $0.0$ $0.3$ FedACG 00:25:53 1.10G
MNIST $0.0$ $0.7$ FedAvg 00:26:08 1.06G
MNIST $0.0$ $0.7$ FedProx 00:17:36 1.07G
MNIST $0.0$ $0.7$ SCAFFOLD 00:17:29 1.03G
MNIST $0.0$ $0.7$ FedACG 00:18:04 1.07G
FMNIST $0.2$ $0.3$ FedAvg 00:25:52 1.10G
FMNIST $0.2$ $0.3$ FedProx 00:26:09 1.14G
FMNIST $0.2$ $0.3$ SCAFFOLD 00:25:58 1.07G
FMNIST $0.2$ $0.3$ FedACG 00:26:18 1.13G
FMNIST $0.2$ $0.7$ FedAvg 00:25:42 1.07G
FMNIST $0.2$ $0.7$ FedProx 00:26:24 1.09G
FMNIST $0.2$ $0.7$ SCAFFOLD 00:26:22 1.06G
FMNIST $0.2$ $0.7$ FedACG 00:25:55 1.10G
FMNIST $0.0$ $0.3$ FedAvg 00:25:50 1.08G
FMNIST $0.0$ $0.3$ FedProx 00:26:28 1.10G
FMNIST $0.0$ $0.3$ SCAFFOLD 00:26:10 1.06G
FMNIST $0.0$ $0.3$ FedACG 00:26:25 1.11G
FMNIST $0.0$ $0.7$ FedAvg 00:25:53 1.07G
FMNIST $0.0$ $0.7$ FedProx 00:26:22 1.10G
FMNIST $0.0$ $0.7$ SCAFFOLD 00:25:38 1.07G
FMNIST $0.0$ $0.7$ FedACG 00:26:21 1.11G
SVHN $0.2$ $0.3$ FedAvg 03:58:13 2.70G
SVHN $0.2$ $0.3$ FedProx 04:14:29 2.74G
SVHN $0.2$ $0.3$ SCAFFOLD 04:11:40 2.59G
SVHN $0.2$ $0.3$ FedACG 04:16:08 2.46G
SVHN $0.2$ $0.7$ FedAvg 04:01:22 2.69G
SVHN $0.2$ $0.7$ FedProx 04:07:41 2.72G
SVHN $0.2$ $0.7$ SCAFFOLD 04:01:41 2.57G
SVHN $0.2$ $0.7$ FedACG 04:14:45 2.46G
SVHN $0.0$ $0.3$ FedAvg 03:53:31 2.69G
SVHN $0.0$ $0.3$ FedProx 04:03:16 2.72G
SVHN $0.0$ $0.3$ SCAFFOLD 03:58:43 2.56G
SVHN $0.0$ $0.3$ FedACG 04:04:32 2.42G
SVHN $0.0$ $0.7$ FedAvg 03:57:38 2.67G
SVHN $0.0$ $0.7$ FedProx 04:10:49 2.73G
SVHN $0.0$ $0.7$ SCAFFOLD 04:00:49 2.57G
SVHN $0.0$ $0.7$ FedACG 04:04:58 2.42G
CIFAR-10 $0.2$ $0.3$ FedAvg 04:48:29 13.67G
CIFAR-10 $0.2$ $0.3$ FedProx 05:22:40 13.74G
CIFAR-10 $0.2$ $0.3$ SCAFFOLD 06:30:16 14.68G
CIFAR-10 $0.2$ $0.3$ FedACG 05:51:08 13.33G
CIFAR-10 $0.2$ $0.7$ FedAvg 04:55:07 13.62G
CIFAR-10 $0.2$ $0.7$ FedProx 05:27:18 13.69G
CIFAR-10 $0.2$ $0.7$ SCAFFOLD 06:30:14 14.54G
CIFAR-10 $0.2$ $0.7$ FedACG 05:50:24 13.34G
CIFAR-10 $0.0$ $0.3$ FedAvg 04:55:51 13.65G
CIFAR-10 $0.0$ $0.3$ FedProx 05:31:49 13.72G
CIFAR-10 $0.0$ $0.3$ SCAFFOLD 06:33:56 14.55G
CIFAR-10 $0.0$ $0.3$ FedACG 05:51:33 13.35G
CIFAR-10 $0.0$ $0.7$ FedAvg 04:59:05 13.64G
CIFAR-10 $0.0$ $0.7$ FedProx 05:18:14 13.72G
CIFAR-10 $0.0$ $0.7$ SCAFFOLD 06:30:40 14.67G
CIFAR-10 $0.0$ $0.7$ FedACG 05:40:13 13.38G
CIFAR-100 $0.2$ $0.3$ FedAvg 05:07:17 13.66G
CIFAR-100 $0.2$ $0.3$ FedProx 05:27:38 13.78G
CIFAR-100 $0.2$ $0.3$ SCAFFOLD 06:34:27 14.61G
CIFAR-100 $0.2$ $0.3$ FedACG 05:45:00 13.22G
CIFAR-100 $0.2$ $0.7$ FedAvg 05:04:29 13.73G
CIFAR-100 $0.2$ $0.7$ FedProx 05:29:20 13.78G
CIFAR-100 $0.2$ $0.7$ SCAFFOLD 06:29:29 14.65G
CIFAR-100 $0.2$ $0.7$ FedACG 05:45:25 13.26G
CIFAR-100 $0.0$ $0.3$ FedAvg 04:53:37 13.71G
CIFAR-100 $0.0$ $0.3$ FedProx 05:25:04 13.82G
CIFAR-100 $0.0$ $0.3$ SCAFFOLD 06:30:22 14.56G
CIFAR-100 $0.0$ $0.3$ FedACG 05:43:58 13.24G
CIFAR-100 $0.0$ $0.7$ FedAvg 04:55:26 13.71G
CIFAR-100 $0.0$ $0.7$ FedProx 05:14:38 13.76G
CIFAR-100 $0.0$ $0.7$ SCAFFOLD 06:21:14 14.69G
CIFAR-100 $0.0$ $0.7$ FedACG 05:38:40 13.25G
TinyImageNet $0.2$ $0.3$ MOON 18:39:30 26.12G
TinyImageNet $0.2$ $0.3$ FedProx 16:22:46 25.01G
TinyImageNet $0.2$ $0.3$ FedACG 17:02:45 23.38G
TinyImageNet $0.2$ $0.7$ MOON 18:08:26 26.20G
TinyImageNet $0.2$ $0.7$ FedProx 15:49:21 25.03G
TinyImageNet $0.2$ $0.7$ FedACG 16:46:22 23.43G
TinyImageNet $0.0$ $0.3$ MOON 18:09:04 26.17G
TinyImageNet $0.0$ $0.3$ FedProx 15:22:17 25.12G
TinyImageNet $0.0$ $0.3$ FedACG 16:31:42 23.34G
TinyImageNet $0.0$ $0.7$ MOON 17:56:09 26.09G
TinyImageNet $0.0$ $0.7$ FedProx 15:59:06 25.02G
TinyImageNet $0.0$ $0.7$ FedACG 16:40:21 23.27G
AG News $0.8$ $0.3$ FedAvg 01:04:40 9.97G
AG News $0.8$ $0.3$ FedProx 01:05:43 10.02G
AG News $0.8$ $0.3$ SCAFFOLD 02:29:40 8.61G
AG News $0.8$ $0.3$ FedACG 01:19:28 8.62G
AG News $0.8$ $0.7$ FedAvg 01:03:21 9.96G
AG News $0.8$ $0.7$ FedProx 01:05:03 9.99G
AG News $0.8$ $0.7$ SCAFFOLD 02:29:45 8.62G
AG News $0.8$ $0.7$ FedACG 01:16:25 8.62G
AG News $0.0$ $0.3$ FedAvg 01:02:22 9.98G
AG News $0.0$ $0.3$ FedProx 01:03:02 9.99G
AG News $0.0$ $0.3$ SCAFFOLD 02:27:12 8.59G
AG News $0.0$ $0.3$ FedACG 01:15:39 8.67G
AG News $0.0$ $0.7$ FedAvg 01:00:31 9.96G
AG News $0.0$ $0.7$ FedProx 01:04:21 10.00G
AG News $0.0$ $0.7$ SCAFFOLD 02:28:45 8.64G
AG News $0.0$ $0.7$ FedACG 01:13:27 8.65G
LABEL:table:resource_100_appendix presents a comprehensive overview of average CPU time and average maximum memory usage for our experiments with 100 clients. The average is taken over three random seeds.
Table 16: Average CPU time and average maximum memory usage across three random seeds for our experiments with 100 clients.
Dataset Cross Alpha Method CPU Time Max Mem
MNIST $0.2$ $0.3$ MOON 00:59:59 1.29G
MNIST $0.2$ $0.3$ FedProx 00:57:34 1.17G
MNIST $0.2$ $0.3$ SCAFFOLD 00:56:08 1.11G
MNIST $0.2$ $0.3$ FedACG 00:57:24 1.16G
MNIST $0.2$ $0.7$ MOON 00:59:29 1.29G
MNIST $0.2$ $0.7$ FedProx 00:57:02 1.17G
MNIST $0.2$ $0.7$ SCAFFOLD 00:55:34 1.13G
MNIST $0.2$ $0.7$ FedACG 00:56:42 1.15G
MNIST $0.0$ $0.3$ MOON 00:59:09 1.28G
MNIST $0.0$ $0.3$ FedProx 00:58:24 1.15G
MNIST $0.0$ $0.3$ SCAFFOLD 00:56:19 1.13G
MNIST $0.0$ $0.3$ FedACG 00:57:24 1.14G
MNIST $0.0$ $0.7$ MOON 01:00:03 1.27G
MNIST $0.0$ $0.7$ FedProx 00:58:13 1.16G
MNIST $0.0$ $0.7$ SCAFFOLD 00:55:47 1.12G
MNIST $0.0$ $0.7$ FedACG 00:57:15 1.17G
FMNIST $0.2$ $0.3$ MOON 00:59:49 1.27G
FMNIST $0.2$ $0.3$ FedProx 00:57:43 1.15G
FMNIST $0.2$ $0.3$ SCAFFOLD 00:55:53 1.13G
FMNIST $0.2$ $0.3$ FedACG 00:57:33 1.16G
FMNIST $0.2$ $0.7$ MOON 00:59:47 1.29G
FMNIST $0.2$ $0.7$ FedProx 00:57:40 1.16G
FMNIST $0.2$ $0.7$ SCAFFOLD 00:56:19 1.11G
FMNIST $0.2$ $0.7$ FedACG 00:57:41 1.16G
FMNIST $0.0$ $0.3$ MOON 00:59:55 1.28G
FMNIST $0.0$ $0.3$ FedProx 00:56:43 1.17G
FMNIST $0.0$ $0.3$ SCAFFOLD 00:56:31 1.13G
FMNIST $0.0$ $0.3$ FedACG 00:57:57 1.17G
FMNIST $0.0$ $0.7$ MOON 00:59:52 1.27G
FMNIST $0.0$ $0.7$ FedProx 00:58:14 1.16G
FMNIST $0.0$ $0.7$ SCAFFOLD 00:56:14 1.12G
FMNIST $0.0$ $0.7$ FedACG 00:57:09 1.15G
SVHN $0.2$ $0.3$ MOON 11:38:02 6.44G
SVHN $0.2$ $0.3$ FedProx 10:38:06 5.99G
SVHN $0.2$ $0.3$ SCAFFOLD 10:29:37 5.85G
SVHN $0.2$ $0.3$ FedACG 10:55:30 5.17G
SVHN $0.2$ $0.7$ MOON 12:11:57 6.42G
SVHN $0.2$ $0.7$ FedProx 11:14:06 6.09G
SVHN $0.2$ $0.7$ SCAFFOLD 11:04:21 5.84G
SVHN $0.2$ $0.7$ FedACG 10:56:34 5.17G
SVHN $0.0$ $0.3$ MOON 11:50:13 6.48G
SVHN $0.0$ $0.3$ FedProx 10:22:12 6.06G
SVHN $0.0$ $0.3$ SCAFFOLD 10:14:38 5.96G
SVHN $0.0$ $0.3$ FedACG 10:51:27 5.23G
SVHN $0.0$ $0.7$ MOON 11:32:17 6.47G
SVHN $0.0$ $0.7$ FedProx 11:03:23 6.09G
SVHN $0.0$ $0.7$ SCAFFOLD 10:50:04 5.77G
SVHN $0.0$ $0.7$ FedACG 10:36:18 5.19G
CIFAR-10 $0.2$ $0.3$ MOON 16:52:19 61.25G
CIFAR-10 $0.2$ $0.3$ FedProx 12:59:18 57.32G
CIFAR-10 $0.2$ $0.3$ SCAFFOLD 20:22:25 56.64G
CIFAR-10 $0.2$ $0.3$ FedACG 14:38:38 53.22G
CIFAR-10 $0.2$ $0.7$ MOON 17:31:36 61.08G
CIFAR-10 $0.2$ $0.7$ FedProx 13:19:24 57.28G
CIFAR-10 $0.2$ $0.7$ SCAFFOLD 20:15:41 56.52G
CIFAR-10 $0.2$ $0.7$ FedACG 14:25:16 53.29G
CIFAR-10 $0.0$ $0.3$ MOON 17:20:51 61.31G
CIFAR-10 $0.0$ $0.3$ FedProx 13:24:55 57.27G
CIFAR-10 $0.0$ $0.3$ SCAFFOLD 20:25:56 56.51G
CIFAR-10 $0.0$ $0.3$ FedACG 14:48:45 53.14G
CIFAR-10 $0.0$ $0.7$ MOON 17:26:10 61.26G
CIFAR-10 $0.0$ $0.7$ FedProx 13:23:31 57.17G
CIFAR-10 $0.0$ $0.7$ SCAFFOLD 20:03:12 56.40G
CIFAR-10 $0.0$ $0.7$ FedACG 14:50:02 53.19G
CIFAR-100 $0.2$ $0.3$ MOON 19:23:54 62.38G
CIFAR-100 $0.2$ $0.3$ FedProx 14:53:36 57.47G
CIFAR-100 $0.2$ $0.3$ SCAFFOLD 21:45:50 58.21G
CIFAR-100 $0.2$ $0.3$ FedACG 16:31:56 54.33G
CIFAR-100 $0.2$ $0.7$ MOON 19:05:52 62.30G
CIFAR-100 $0.2$ $0.7$ FedProx 14:58:14 57.43G
CIFAR-100 $0.2$ $0.7$ SCAFFOLD 21:39:22 58.15G
CIFAR-100 $0.2$ $0.7$ FedACG 16:18:23 54.22G
CIFAR-100 $0.0$ $0.3$ MOON 17:36:12 62.01G
CIFAR-100 $0.0$ $0.3$ FedProx 13:27:23 57.11G
CIFAR-100 $0.0$ $0.3$ SCAFFOLD 20:51:50 58.11G
CIFAR-100 $0.0$ $0.3$ FedACG 15:04:23 54.19G
CIFAR-100 $0.0$ $0.7$ MOON 17:43:03 62.00G
CIFAR-100 $0.0$ $0.7$ FedProx 13:26:15 57.19G
CIFAR-100 $0.0$ $0.7$ SCAFFOLD 20:09:42 58.26G
CIFAR-100 $0.0$ $0.7$ FedACG 14:50:34 54.36G
AG News $0.8$ $0.3$ MOON 07:28:40 39.48G
AG News $0.8$ $0.3$ FedProx 05:05:26 42.81G
AG News $0.8$ $0.3$ SCAFFOLD 10:45:28 31.41G
AG News $0.8$ $0.3$ FedACG 06:12:14 35.95G
AG News $0.8$ $0.7$ MOON 07:26:40 39.39G
AG News $0.8$ $0.7$ FedProx 05:14:55 42.77G
AG News $0.8$ $0.7$ SCAFFOLD 11:04:57 31.39G
AG News $0.8$ $0.7$ FedACG 06:02:33 35.95G
AG News $0.0$ $0.3$ MOON 07:28:01 39.42G
AG News $0.0$ $0.3$ FedProx 04:59:05 42.79G
AG News $0.0$ $0.3$ SCAFFOLD 11:25:17 33.66G
AG News $0.0$ $0.3$ FedACG 05:49:40 35.98G
AG News $0.0$ $0.7$ MOON 07:29:04 39.55G
AG News $0.0$ $0.7$ FedProx 05:11:57 42.84G
AG News $0.0$ $0.7$ SCAFFOLD 10:57:00 31.42G
AG News $0.0$ $0.7$ FedACG 05:51:58 36.01G
Tinyimagenet $0.2$ $0.3$ MOON 71:04:06 107.02G
Tinyimagenet $0.2$ $0.3$ FedProx 60:05:53 104.66G
Tinyimagenet $0.2$ $0.3$ SCAFFOLD 72:56:11 95.09G
Tinyimagenet $0.2$ $0.3$ FedACG 63:35:32 92.64G
Tinyimagenet $0.2$ $0.7$ MOON 77:32:22 106.88G
Tinyimagenet $0.2$ $0.7$ FedProx 64:33:53 104.34G
Tinyimagenet $0.2$ $0.7$ SCAFFOLD 79:16:25 94.82G
Tinyimagenet $0.2$ $0.7$ FedACG 66:31:58 92.48G
Tinyimagenet $0.0$ $0.3$ MOON 70:13:51 106.77G
Tinyimagenet $0.0$ $0.3$ FedProx 58:26:32 104.38G
Tinyimagenet $0.0$ $0.3$ SCAFFOLD 73:27:39 95.02G
Tinyimagenet $0.0$ $0.3$ FedACG 61:50:21 92.33G
Tinyimagenet $0.0$ $0.7$ MOON 69:09:45 106.72G
Tinyimagenet $0.0$ $0.7$ FedProx 59:04:02 104.18G
Tinyimagenet $0.0$ $0.7$ SCAFFOLD 71:57:32 95.24G
Tinyimagenet $0.0$ $0.7$ FedACG 61:02:40 92.33G
Appendix E Proof of Theorem 1
Based on the local training rule and global weighted average aggregation rule, the different between two global models is
$\bm{w}^{(g+1)}-\bm{w}^{(g)}=\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\nabla\overline{F}_{k}^{(g)}$ (6)
where $\nabla\overline{F}_{k}^{(g)}$ is the accumulated gradient across $e_{k}^{(g)}$ local steps given by
$\nabla\overline{F}_{k}^{(g)}=\left(\bm{w}_{k}^{(g),e_{k}^{(g)}}-\bm{w}^{(g)}\right)=-\eta^{(g)}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g)},d\right)}{B_{k}^{(g)}}$ (7)
Taking the expectation of (7) and using the fact that mini-batch sampling is conducted uniformly at random without replacement, we have
$\mathbb{E}\left[\nabla\overline{F}_{k}^{(g)}\right]=-\eta^{(g)}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f\left(\bm{w}_{k}^{(g)},d\right)}{D_{k}^{(g)}}$ (8)
Using the $\beta$-smoothness of the global loss function (Assumption 1), we have
$F^{(g)}(\bm{w}^{(g+1)})\leq F^{(g)}(\bm{w}^{(g)})+\nabla F^{(g)}(\bm{w}^{(g)})^{\top}(\bm{w}^{(g+1)}-\bm{w}^{(g)})+\frac{\beta}{2}\|\bm{w}^{(g+1)}-\bm{w}^{(g)}\|^{2}.$ (9)
Plugging (6) and (7) into (9) and taking the expectation give
$\mathbb{E}_{g}\left[F^{(g)}(\bm{w}^{(g+1)})\right]$ (10)
$\leq F^{(g)}(\bm{w}^{(g)})-\eta^{(g)}\nabla F^{(g)}(\bm{w}^{(g)})^{\top}\mathbb{E}_{g}\left[\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g),e-1},d\right)}{D_{k}^{(g)}}\right]$
$\quad+\frac{\beta}{2}\mathbb{E}_{g}\left[\left\|\eta^{(g)}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g),e-1},d\right)}{B_{k}^{(g)}}\right\|^{2}\right]$
Using the fact that for any two real-valued vectors $\bm{a}$ and $\bm{b}$ with the same length, we have: $2\bm{a}^{\top}\bm{b}=\|\bm{a}\|^{2}+\|\bm{b}\|^{2}-\|\bm{a}-\bm{b}\|^{2},$, we further obtain
$\mathbb{E}_{g}\left[F^{(g)}(\bm{w}^{(g+1)})\right]\leq F^{(g)}(\bm{w}^{(g)})-\frac{\eta^{(g)}}{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\left\|\nabla F^{(g)}(\bm{w}^{(g)})\right\|^{2}$
$\quad+\left\|\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g),e-1},d\right)}{D_{k}^{(g)}}\right\|^{2}$
$\quad+\left\|\nabla F^{(g)}(\bm{w}^{(g)})-\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g),e-1},d\right)}{D_{k}^{(g)}}\right\|^{2}\Bigg{]}$
$\quad+\frac{\beta(\eta^{(g)})^{2}}{2}\underbrace{\mathbb{E}_{g}\Bigg{[}\left\|\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g),e-1},d\right)}{B_{k}^{(g)}}\right\|^{2}\Bigg{]}}_{\text{(a)}}$ (11)
We first bound term (a) as follows.
$\mathbb{E}_{g}\left[\left\|\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}\right\|^{2}\right]$
$=\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\Bigg{(}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}-\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}$
$\quad+\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{)}\Bigg{\|}^{2}\Bigg{]}$
$\overset{(i)}{\leq}2\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\Bigg{(}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}-\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{)}\Bigg{\|}^{2}\Bigg{]}$
$\quad+2\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}$
$\overset{(ii)}{\leq}2\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}-\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}\Bigg{]}$
$\quad+2\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}$
$\overset{(iii)}{\leq}2\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{(\sigma_{k}^{(g),e-1})^{2}}{B_{k}^{(g)}}$ (12)
$\quad+2\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}.$
where (i) we have used the Cauchy-Schwarz inequality $\|\bm{a}+\bm{b}\|^{2}\leq 2\|\bm{a}\|^{2}+2\|\bm{b}\|^{2}$, (ii) we used the convex property of 2-norm, (iii) we used Eq (3.5) in chapter 3 of [51] and the definition of variance of the gradients $(\sigma_{k}^{(g),e-1})^{2}$ evaluated at the particular local step $e-1$ for the parameter realization $\bm{w}_{k}^{(g),e-1}$. Expanding $(\sigma_{k}^{(g),e-1})^{2}$ yields
$(\sigma_{k}^{(g),e-1})^{2}$
$=\frac{\sum_{d\in\mathcal{D}_{k}^{(g)}}\left\|\nabla f(\bm{w}_{k}^{(g),e-1},d)-\frac{\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\nabla f(\bm{w}_{k}^{(g),e-1},d^{\prime})}{D_{k}^{(g)}}\right\|^{2}}{D_{k}^{(g)}-1}$
$=\frac{\sum_{d\in\mathcal{D}_{k}^{(g)}}\left(\frac{1}{D_{k}^{(g)}}\right)^{2}\left\|D_{k}^{(g)}\nabla f(\bm{w}_{k}^{(g),e-1},d)-\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\nabla f(\bm{w}_{k}^{(g),e-1},d^{\prime})\right\|^{2}}{D_{k}^{(g)}-1}$
$\overset{(i)}{\leq}\frac{\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{(D_{k}^{(g)}-1)}{(D_{k}^{(g)})^{2}}\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\left\|\nabla f(\bm{w}_{k}^{(g),e-1},d)-\nabla f(\bm{w}_{k}^{(g),e-1},d^{\prime})\right\|^{2}}{D_{k}^{(g)}-1}$
$\leq\frac{1}{(D_{k}^{(g)})^{2}}(\Theta^{(g)})^{2}\sum_{d\in\mathcal{D}_{k}^{(g)}}\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\left\|d-d^{\prime}\right\|^{2}$
$\leq\frac{1}{(D_{k}^{(g)})^{2}}(\Theta^{(g)})^{2}\sum_{d\in\mathcal{D}_{k}^{(g)}}\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\left\|d-d^{\prime}\right\|^{2}$
$\leq\frac{(\Theta^{(g)})^{2}}{(D_{k}^{(g)})^{2}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\left\|d-\bm{\lambda}_{k}^{(g)}-d^{\prime}+\bm{\lambda}_{k}^{(g)}\right\|^{2}$
$\leq\frac{(\Theta^{(g)})^{2}}{(D_{k}^{(g)})^{2}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\Big{(}\left\|d-\bm{\lambda}_{k}^{(g)}\right\|^{2}+\left\|d^{\prime}-\bm{\lambda}_{k}^{(g)}\right\|^{2}+2(d-\bm{\lambda}_{k}^{(g)})^{\top}(d^{\prime}-\bm{\lambda}_{k}^{(g)})\Big{)}$
$\overset{(ii)}{\leq}\frac{(\Theta^{(g)})^{2}}{(D_{k}^{(g)})^{2}}\Bigg{(}D_{k}^{(g)}\sum_{d\in\mathcal{D}_{k}^{(g)}}\left\|d-\bm{\lambda}_{k}^{(g)}\right\|^{2}+D_{k}^{(g)}\sum_{d^{\prime}\in\mathcal{D}_{k}^{(g)}}\left\|d^{\prime}-\bm{\lambda}_{k}^{(g)}\right\|^{2}\Bigg{)}$
$\leq\frac{2(\Theta^{(g)})^{2}}{D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}.$ (13)
where $d$ denotes the feature vector of data point $d$, and $\bm{\lambda}_{k}^{(g)}$ and $(\widetilde{\sigma}_{k}^{(g)})^{2}$ denote the mean and sample variance of data points in $\mathcal{D}_{k}^{(g)}$, which are gradient independent. Also, in (i) of (E), we used the Cauchy-Schwarz inequality, and in (ii), we used the fact that $\sum_{d\in\mathcal{D}_{k}^{(g)}}\left(d-\bm{\lambda}_{k}^{(g)}\right)=0.$ Plugging (E) into (E) gives
$\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g),e}}\Bigg{\|}^{2}\Bigg{]}$
$\leq 4(\Theta^{(g)})^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}\sum_{e=1}^{e_{k}^{(g)}}\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{1}{B_{k}^{(g)}}$
$+2\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}.$
Using this, (E) becomes
$\mathbb{E}_{g}\left[F^{(g)}(\bm{w}^{(g+1)})\right]\leq F^{(g)}(\bm{w}^{(g)})-\frac{\eta^{(g)}}{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\left\|\nabla F^{(g)}(\bm{w}^{(g)})\right\|^{2}$
$\quad+\left\|\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g),e-1},d\right)}{D_{k}^{(g)}}\right\|^{2}$
$\quad+\left\|\nabla F^{(g)}(\bm{w}^{(g)})-\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g),e}}\frac{\nabla f\left(\bm{w}_{k}^{(g),e-1},d\right)}{D_{k}^{(g)}}\right\|^{2}\Bigg{]}$
$\quad+2\beta(\eta^{(g)})^{2}(\Theta^{(g)})^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}\sum_{e=1}^{e_{k}^{(g)}}\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{1}{B_{k}^{(g)}}$
$+\beta(\eta^{(g)})^{2}\Bigg{\|}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}.$ (14)
Rearranging the terms, we have
$\mathbb{E}_{g}\big{[}F^{(g)}(\bm{w}^{(g+1)})\big{]}$
$\leq F^{(g)}(\bm{w}^{(g)})-\frac{\eta^{(g)}}{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\big{[}\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}\big{]}$
$\quad+\Big{(}\beta(\eta^{(g)})^{2}-\frac{\eta^{(g)}}{2}\Big{)}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}\Bigg{]}$
$\quad+\underbrace{\frac{\eta^{(g)}}{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F^{(g)}(\bm{w}^{(g)})-\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}\Bigg{]}}_{\text{(b)}}$
$\quad+2\beta(\Theta^{(g)})^{2}(\eta^{(g)})^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{(\widetilde{\sigma}_{k}^{(g)})^{2}}{D_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{1}{B_{k}^{(g)}}.$ (15)
Term (b) in (E) can be upper bounded as
$\frac{\eta^{(g)}}{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F^{(g)}(\bm{w}^{(g)})-\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}\Bigg{]}$
$=\frac{\eta^{(g)}}{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F^{(g)}(\bm{w}^{(g)})-\frac{1}{e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}$
$\quad+\left(\frac{1}{e_{k}^{(g)}}-1\right)\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}\Bigg{]}$
$\leq\eta^{(g)}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\underbrace{\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F^{(g)}(\bm{w}^{(g)})-\frac{1}{e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\nabla f(\bm{w}_{k}^{(g),e-1},d)\Bigg{\|}^{2}\Bigg{]}}_{\text{(b1)}}$
$\quad+\eta^{(g)}\Bigg{(}\frac{1}{e_{\min}^{(g)}}-1\Bigg{)}^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\nabla f(\bm{w}_{k}^{(g),e-1},d)\Bigg{\|}^{2}\Bigg{]}.$ (16)
where $e_{\text{min}}^{(g)}=\min_{k\in\mathcal{K}^{(g)}}e_{k}^{(g)}$ is the mininum number of local steps in the $g$-th round. Term (b1) in (16) can be upper bounded as
$\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F^{(g)}(\bm{w}^{(g)})-\frac{1}{e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}\Bigg{]}$
$=\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F^{(g)}(\bm{w}^{(g)})-\frac{1}{e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),e-1})\Bigg{\|}^{2}\Bigg{]}$
$\leq\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F_{k}^{(g)}(\bm{w}^{(g)})-\frac{1}{e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),e-1})\Bigg{\|}^{2}\Bigg{]}$
$\leq\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\Bigg{[}\Bigg{\|}\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g)})-\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),e-1})\Bigg{\|}^{2}\Bigg{]}$
$\leq\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\beta^{2}\mathbb{E}_{g}\Big{[}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}\Big{]}.$ (17)
Using the local update rule, $\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$ can be further upper bounded as
$\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
$=(\eta^{(g)})^{2}\mathbb{E}_{g}\Bigg{\|}\sum_{e^{\prime}=1}^{e-1}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}\Bigg{\|}^{2}$
$=(\eta^{(g)})^{2}\mathbb{E}_{g}\Bigg{\|}\sum_{e^{\prime}=1}^{e-1}\left(\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}-\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\right.$
$\quad\left.+\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\right)\Bigg{\|}^{2}$
$\leq 2(\eta^{(g)})^{2}\mathbb{E}_{g}\Bigg{\|}\sum_{e^{\prime}=1}^{e-1}\left(\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}-\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\right)\Bigg{\|}^{2}$
$\quad+2(\eta^{(g)})^{2}\mathbb{E}_{g}\Bigg{\|}\sum_{e^{\prime}=1}^{e-1}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}$
$\leq 2(\eta^{(g)})^{2}\sum_{e^{\prime}=1}^{e-1}\mathbb{E}_{g}\Bigg{\|}\sum_{d\in\mathcal{B}_{k}^{(g),e}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{B_{k}^{(g)}}-\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}$
$\quad+2(\eta^{(g)})^{2}\mathbb{E}_{g}\Bigg{\|}\sum_{e^{\prime}=1}^{e-1}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}$
$\leq 4(\eta^{(g)})^{2}\sum_{e^{\prime}=1}^{e-1}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+2(\eta^{(g)})^{2}\mathbb{E}_{g}\Bigg{\|}\sum_{e^{\prime}=1}^{e-1}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}.$ (18)
The last term in (18) can be further upper bounded
$2(\eta^{(g)})^{2}\mathbb{E}_{g}\Bigg{\|}\sum_{e^{\prime}=1}^{e-1}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e^{\prime}-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}$
$\leq 2(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\mathbb{E}_{g}\Bigg{\|}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e^{\prime}-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}$
$\leq 2(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\mathbb{E}_{g}\Bigg{\|}\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),e^{\prime}-1})-\nabla F_{k}^{(g)}(\bm{w}^{(g)})+\nabla F_{k}^{(g)}(\bm{w}^{(g)})\Bigg{\|}^{2}$
$\leq 4(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\mathbb{E}_{g}\Bigg{\|}\nabla F_{k}^{(g)}(\bm{w}_{k}^{(g),e^{\prime}-1})-\nabla F_{k}^{(g)}(\bm{w}^{(g)})\Bigg{\|}^{2}$
$\quad+4(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\Bigg{\|}\nabla F_{k}^{(g)}(\bm{w}^{(g)})\Bigg{\|}^{2}$
$\leq 4\beta^{2}(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\mathbb{E}_{g}\Big{\|}\bm{w}_{k}^{(g),e^{\prime}-1}-\bm{w}^{(g)}\Big{\|}^{2}$
$\quad+4(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\Big{\|}\nabla F_{k}^{(g)}(\bm{w}^{(g)})\Big{\|}^{2}.$ (19)
Combining (19) with (18) gives
$\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
$\quad\leq 4(\eta^{(g)})^{2}\sum_{e^{\prime}=1}^{e-1}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+4\beta^{2}(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\mathbb{E}_{g}\|\bm{w}_{k}^{(g),e^{\prime}-1}-\bm{w}^{(g)}\|^{2}$
$\quad+4(\eta^{(g)})^{2}(e-1)\sum_{e^{\prime}=1}^{e-1}\|\nabla F_{k}^{(g)}(\bm{w}^{(g)})\|^{2}.$ (20)
Taking the summation of both sides in (20) across local training steps
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
$\leq 4(\eta^{(g)})^{2}\sum_{e=1}^{e_{k}^{(g)}}\sum_{e^{\prime}=1}^{e-1}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+4\beta^{2}(\eta^{(g)})^{2}\sum_{e=1}^{e_{k}^{(g)}}(e-1)\sum_{e^{\prime}=1}^{e-1}\mathbb{E}_{g}\|\bm{w}_{k}^{(g),e^{\prime}-1}-\bm{w}^{(g)}\|^{2}$
$\quad+4(\eta^{(g)})^{2}\sum_{e=1}^{e_{k}^{(g)}}(e-1)\sum_{e^{\prime}=1}^{e-1}\|\nabla F_{k}^{(g)}(\bm{w}^{(g)})\|^{2}.$
After some simplification, we obtain
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
$\leq 2(\eta^{(g)})^{2}(e_{k}^{(g)}-1)e_{k}^{(g)}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+4\beta^{2}(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
$\quad+4(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)\sum_{e=1}^{e_{k}^{(g)}}\|\nabla F_{k}^{(g)}(\bm{w}^{(g)})\|^{2}.$ (21)
Assuming
$\eta^{(g)}\leq\Big{(}2\beta^{2}\sqrt{e_{k}^{(g)}(e_{k}^{(g)}-1)}\Big{)}^{-1},$
Performing some algebraic operations and combining the same terms on (21) gives
$\sum_{e=1}^{e_{k}^{(g)}}\mathbb{E}_{g}\|\bm{w}^{(g)}-\bm{w}_{k}^{(g),e-1}\|^{2}$
$\leq\frac{2(\eta^{(g)})^{2}(e_{k}^{(g)}-1)e_{k}^{(g)}}{1-4\beta^{2}(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+\frac{4(\eta^{(g)})^{2}(e_{k}^{(g)}(e_{k}^{(g)}-1))}{1-4\beta^{2}(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)}\sum_{e=1}^{e_{k}^{(g)}}\|\nabla F_{k}^{(g)}(\bm{w}^{(g)})\|^{2}.$ (22)
Let $e_{\text{max}}^{(g)}=\max_{k\in\mathcal{K}^{(g)}}e_{k}^{(g)}$ be the maximum number of local steps in the $g$-th round, plugging (22) in (17) gives
$\mathbb{E}_{g}\|\nabla F^{(g)}(\bm{w}^{(g)})-\frac{1}{e_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\nabla f(\bm{w}_{k}^{(g),e-1},d)\|^{2}$
$\leq 2(\eta^{(g)})^{2}\beta^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\frac{e_{k}^{(g)}-1}{1-4\beta^{2}(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+\frac{4\beta^{2}(\eta^{(g)})^{2}(e_{\max}^{(g)}(e_{\max}^{(g)}-1))}{1-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\|\nabla F_{k}^{(g)}(\bm{w}^{(g)})\|^{2}$
$\overset{(i)}{\leq}2(\eta^{(g)})^{2}\beta^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\frac{e_{k}^{(g)}-1}{1-4\beta^{2}(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+\frac{4\beta^{2}(\eta^{(g)})^{2}(e_{\max}^{(g)}(e_{\max}^{(g)}-1))}{1-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)}\left(\zeta_{1}\mathbb{E}_{g}\left\|\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\nabla F_{k}^{(g)}(\bm{w}^{(g)})\right\|^{2}+\zeta_{2}\right)$
$\leq 2(\eta^{(g)})^{2}\beta^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\frac{e_{k}^{(g)}-1}{1-4\beta^{2}(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+\frac{4\beta^{2}(\eta^{(g)})^{2}\big{(}e_{\max}^{(g)}\big{)}^{2}\big{(}e_{\max}^{(g)}-1\big{)}}{1-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)}\Big{(}\zeta_{1}\mathbb{E}_{g}\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}+\zeta_{2}\Big{)}.$ (23)
Plugging (23) into (E) gives
$\mathbb{E}_{g}\big{[}F^{(g)}(\bm{w}^{(g+1)})\big{]}\leq F^{(g)}(\bm{w}^{(g)})+2\beta(\Theta^{(g)})^{2}(\eta^{(g)})^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{(\widetilde{\sigma}_{k}^{(g)})^{2}}{D_{k}^{(g)}}\sum_{e=1}^{e_{k}^{(g)}}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{1}{B_{k}^{(g)}}$
$\quad+(\eta^{(g)})^{3}\beta^{2}\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\frac{e_{k}^{(g)}-1}{1-4\beta^{2}(\eta^{(g)})^{2}e_{k}^{(g)}(e_{k}^{(g)}-1)}\Bigg{(}1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\Bigg{)}\frac{(\Theta^{(g)})^{2}}{B_{k}^{(g)}D_{k}^{(g)}}(\widetilde{\sigma}_{k}^{(g)})^{2}$
$\quad+\underbrace{\left(\beta(\eta^{(g)})^{2}-\frac{\eta^{(g)}}{2}+\eta^{(g)}\Bigg{(}\frac{1}{e_{\min}^{(g)}}-1\Bigg{)}^{2}\right)\sum_{k\in\mathcal{K}^{(g)}}\frac{D_{k}^{(g)}}{D^{(g)}}\mathbb{E}_{g}\Bigg{\|}\sum_{e=1}^{e_{k}^{(g)}}\sum_{d\in\mathcal{D}_{k}^{(g)}}\frac{\nabla f(\bm{w}_{k}^{(g),e-1},d)}{D_{k}^{(g)}}\Bigg{\|}^{2}}_{\text{(c)}}$
$\quad+\frac{2\beta^{2}(\eta^{(g)})^{3}\big{(}e_{\max}^{(g)}\big{)}^{2}\big{(}e_{\max}^{(g)}-1\big{)}}{1-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)}\zeta_{2}$
$\quad+\frac{\eta^{(g)}}{2}\Bigg{(}\frac{4\beta^{2}(\eta^{(g)})^{2}\big{(}e_{\max}^{(g)}\big{)}^{2}\big{(}e_{\max}^{(g)}-1\big{)}}{1-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)}\zeta_{1}-1\Bigg{)}\Bigg{\|}\nabla F^{(g)}(\bm{w}^{(g)})\Bigg{\|}^{2}.$ (24)
To eliminate term (c) in (24), we need to make
$\beta(\eta^{(g)})^{2}-\frac{\eta^{(g)}}{2}+\eta^{(g)}\left(\frac{1}{e_{\min}^{(g)}}-1\right)^{2}<0,$ (25)
we require
$\eta^{(g)}<\left(\frac{1}{2}-\left(\frac{1}{e_{\min}^{(g)}}-1\right)^{2}\right)\frac{1}{\beta}.$ (26)
Also, we aim to make
$\frac{4\beta^{2}(\eta^{(g)})^{2}\big{(}e_{\max}^{(g)}\big{)}^{2}\big{(}e_{\max}^{(g)}-1\big{)}}{1-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)}-1<0.$ (27)
$01-4\beta^{2}(\eta^{(g)})^{2}e_{\max}^{(g)}(e_{\max}^{(g)}-1)>$ 0 . (28)
Thus, we get the condition:
$\eta^{(g)}<\frac{1}{2\beta\sqrt{e_{\max}^{(g)}\big{(}e_{\max}^{(g)}-1\big{)}}}.$ (29)
Assume there exists a set of constants $\Lambda^{(g)}$, $\forall g$, where:
$\frac{4(\eta^{(g)})^{2}\beta^{2}(\Theta^{(g)})^{2}(e_{\text{max}}^{(g)}(e_{\text{max}}^{(g)}-1))}{1-4(\eta^{(g)})^{2}\beta^{2}e_{\text{max}}^{(g)}(e_{\text{max}}^{(g)}-1)}\zeta_{1}$ $<\Lambda^{(g)}<1,\quad\Lambda^{(g)}<\zeta_{1}$ (30)
Then we will have
$\eta^{(g)}$ $\leq\frac{\sqrt{\Lambda^{(g)}}}{2\beta\sqrt{\left(\zeta_{1}+\Lambda^{(g)}\right)e_{\text{max}}^{(g)}\left(e_{\text{max}}^{(g)}-1\right)}}$ (31)
and
$\frac{1}{\left(1-4(\eta^{(g)})^{2}\beta^{2}e_{\text{max}}^{(g)}(e_{\text{max}}^{(g)}-1)\right)}\leq\frac{\zeta_{1}+\Lambda^{(g)}}{\zeta_{1}}$ (32)
Note that the condition in (31) is stricter than (29). Putting altogether, we get
$\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}$
$\leq\frac{2\left(F^{(g)}(\bm{w}^{(g)})-\mathbb{E}_{g}\left[F^{(g)}(\bm{w}^{(g+1)})\right]\right)}{\eta^{(g)}(1-\Lambda^{(g)})}$
$\quad+\frac{4\beta(\Theta^{(g)})^{2}\eta^{(g)})}{1-\Lambda^{(g)}}\sum_{k\in\mathcal{K}^{(g)}}\frac{(\widetilde{\sigma}_{k}^{(g)})^{2}}{D_{k}^{(g)}}e_{k}^{(g)}\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{1}{B_{k}^{(g)}}$
$\quad+\frac{2\beta^{2}(\eta^{(g)})^{2}}{D^{(g)}(1-\Lambda^{(g)})}\sum_{k\in\mathcal{K}^{(g)}}(e_{k}^{(g)}-1)\frac{\zeta_{1}+\Lambda^{(g)}}{\zeta_{1}}\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{(\Theta^{(g)})(\widetilde{\sigma}_{k}^{(g)})^{2}}{B_{k}^{(g)}}$
$\quad+4\beta^{2}(\eta^{(g)})^{2}(e_{\text{max}}^{(g)})(e_{\text{max}}^{(g)}-1)\frac{\zeta_{2}}{1-\Lambda^{(g)}}\frac{(\zeta_{1}+\Lambda^{(g)})}{\zeta_{1}}$
Summing from 1 to G and dividing by G and applying $\frac{\zeta_{1}+\Lambda^{(g)}}{\zeta_{1}}<2$ lead to
$\frac{1}{G}\sum_{g=1}^{G}\|\nabla F^{(g)}(\bm{w}^{(g)})\|^{2}$
$\leq\frac{1}{G}\sum_{g=1}^{G}\frac{2\left(F^{(g)}(\bm{w}^{(g)})-\mathbb{E}_{g}\left[F^{(g)}(\bm{w}^{(g+1)})\right]\right)}{\eta^{(g)}(1-\Lambda^{(g)})}$
$\quad+\frac{1}{G}\sum_{g=1}^{G}\frac{4\beta(\Theta^{(g)})^{2}(\eta^{(g)})}{1-\Lambda^{(g)}}\sum_{k\in\mathcal{K}^{(g)}}\frac{(\widetilde{\sigma}_{k}^{(g)})^{2}}{D_{k}^{(g)}}e_{k}^{(g)}\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{1}{B_{k}^{(g)}}$
$\quad+\frac{1}{G}\sum_{g=1}^{G}\frac{4\beta^{2}(\eta^{(g)})^{2}}{D^{(g)}(1-\Lambda^{(g)})}\sum_{k\in\mathcal{K}^{(g)}}(e_{k}^{(g)}-1)\left(1-\frac{B_{k}^{(g)}}{D_{k}^{(g)}}\right)\frac{(\Theta^{(g)})(\widetilde{\sigma}_{k}^{(g)})^{2}}{B_{k}^{(g)}}$
$\quad+\frac{1}{G}\sum_{g=1}^{G}8\beta^{2}(\eta^{(g)})^{2}(e_{\text{max}}^{(g)})(e_{\text{max}}^{(g)}-1)\frac{\zeta_{2}}{1-\Lambda^{(g)}}$